### v1

# EVA-Score: Evaluation of Long-form Summarization on Informativeness through Extraction and Validation

Abstract
Summarization is a fundamental task in natural language processing (NLP) and since large language models (LLMs), such as GPT-4 and Claude, come out, increasing attention has been paid to long-form summarization whose input sequences are much longer, indicating more information contained. The current evaluation metrics either use similarity-based metrics like ROUGE and BERTScore which rely on similarity and fail to consider informativeness or LLM-based metrics, lacking quantitative analysis of information richness and are rather subjective. In this paper, we propose a new evaluation metric called EVA-Score using Atomic Fact Chain Generation and Document-level Relation Extraction together to automatically calculate the informativeness and give a definite number as an information score. Experiment results show that our metric shows a state-of-the-art correlation with humans. We also re-evaluate the performance of LLMs on long-form summarization comprehensively from the information aspect, forecasting future ways to use LLMs for long-form summarization.
EVA-Score: Evaluation of Long-form Summarization on Informativeness through Extraction and Validation
Yuchen Fan1thanks:  The work was done when he worked as an intern at Tsinghua University, China., Xin Zhong1, Chengsi Wang1, Gaochen Wu2, Bowen Zhou2thanks:  Corresponding authors 1 Beijing University of Posts and Telecommunications 2 Department of Electronic Engineering, Tsinghua University yuchenfan48@gmail.com zhoubowen@tsinghua.edu.cn
1 Introduction
Figure 1: Examples show the dilemma of ROUGE. The blue annotation is the atomic information contained in the text. Even though Candidate Summarization 1 is rewarded with a higher ROUGE Score, it contains less information than Candidate Summarization 2. It inspires us that ROUGE and BERTScore are not reasonable for evaluating summarization.
Summarization has long been a fundamental task in NLP and since LLMs came out (OpenAI et al., 2024; Touvron et al., 2023), many research leverages LLMs for summarization (Zhang et al., 2023b; Wang et al., 2023b). More recently, works on summarization moved to Long-form scenarios where the input sequence length is extremely larger, resulting in more information contained in it. Traditional evaluation metrics may be not suitable for evaluation of long-form summarization due to the great informativeness, thus calling for new evaluation metrics.
Past evaluation metrics for summarization include some static evaluation methods, like ROUGE-k and ROUGE-l (Lin, 2004) which calculate the k-gram overlap or the longest common substring between the reference and the candidate. With the development of pre-trained language models, BERTScore (Zhang et al., 2020) is widely used as a new evaluation metric based on pre-trained word embeddings. More recently, some works have begun to pay attention to evaluating long-form summarization (Kim et al., 2024; Chang et al., 2024) which uses prompts to define criteria to assign different scores for different aspects, like fluency, consistency, and coherence. However, traditional and LLM-based evaluation metrics face their dilemmas respectively.
Figure 2: Pipeline of our evaluation metric. We begin with atomic fact generation. For generated facts, we use LLMs to check whether they can be treated as a logic chain and use Document-level Relation Extraction as a complement. Finally, when analyzing the truthfulness of a given fact, we use retrieval to find several reference facts and provide previous states.
For Similarity-based Metrics Since they are based on similarity, they cannot represent informativeness and are unsuitable in such an LLM era, where responses from LLMs are flexible and contain many synonyms and different syntax. For LLM-based Metrics They are insensitive to nuanced differences existing in summaries and their quantitative results are not reliable even if given detailed criteria for evaluation due to the black-box property.
We propose a new evaluation metric whose result is an objective and quantitative analysis but can fully catch the information contained in the summary. We call this metric Extraction-and-Validation Score. We begin with extracting atomic facts from reference and candidate summaries. Different from FActScore (Min et al., 2023), we claim that some information cannot live without a context, so we treat them as part of logic chains. Afterward, we introduce Document-level Relation Extraction to remedy the loss of hierarchical relations. Then for each atomic fact to be verified, we use information retrieval to find the most relevant atomic facts as references and prompt LLMs to focus only on newly added information and decide whether it is entailed or not. Intuitively, LLMs for evaluation can be more accurate in single verification than in complex situations. We ensure that only one piece of information is evaluated at a time, simplifying the validation process. And by introducing Document-level Relations, we form a more comprehensive information set, maintaining the completeness of the evaluation.
We conduct comprehensive experiments by comparing EVA-Score with current evaluation metrics, including traditional and LLM-based. The results show that EVA-Score achieves the best human alignment compared to SOTA past evaluation metric, BERTScore by 0.166, 0.120, and 0.09 in Pearson, Spearman, and Kendall correlation respectively at the text level. At the system level, it shows human-comparable evaluation, with 0.975 in Pearson correlation and 1.0 in both Spearman and Kendall correlation. We also re-evaluate long-form summarization based on EVA-Score and the experiment results show the informativeness of different LLMs on long-form summarization. The results show the great performance of GPT-4 when the context becomes longer, ranking first in arXiv and Gov Report and second in Pubmed and BookSum. It also shows the promising performance of Mixtral-7B-Instruct for long-form summarization.
Our main contributions are as follows:
  • We propose a new evaluation metric that can generate high-quality evaluation results of a given summary, based on atomic fact chain generation and document-level relation extraction, by providing a numerical score and steps to obtain this score.
  • We conduct comprehensive experiments on several long-form summarization datasets, with the number of tokens of each record ranging from 1k to 14k. Experiment results show that our evaluation metric shows superiority in every correlation with humans.
  • We re-evaluate long-form summarization concerning informativeness which is the first to evaluate from the information aspect as far as we know. In our experiments, the state-of-the-art model GPT-4 still performs better than any other models in long context settings but lags behind some assistant models in short context settings.
2 Related Work
2.1 Long-form Summarization
Long-form summarization requires language models to understand the purpose of the context and capture core information from a long context which is a struggle for language models. During the last few years, much work from different fields regarding long-form summarization has been studied such as news reports (See et al., 2017), government reports (Cohan et al., 2018), academic papers and books (Kryściński et al., 2022), posing great challenges for language models (Lewis et al., 2019; Raffel et al., 2023). Since ChatGPT and GPT-4 (OpenAI et al., 2024) come out, LLMs show great potential to deal with long-form summarization (Lu et al., 2023; Zhang et al., 2023a; Wang et al., 2023a) because of their huge amount of pre-training corpus. However, when the context length becomes much larger, LLMs still have trouble because of limitations for input sequences and knowledge forgetting when reading a long context. More recently, some research aims at exploring LLMs which can handle longer context paving the way for long-form summarization. (Su et al., 2023; Li* et al., 2023) Yet, summarizing report-level or book-level texts is still a difficult task. (Chang et al., 2024)
2.2 Evaluation of Summarization
Tradition evaluation metrics for summarization include ROUGE variants, like ROUGE-k and ROUGE-L and BERTScore (Zhang et al., 2020). However, they only reflect the similarity between candidates and references without the information contained in the prediction which is not enough for us to choose a better summary or quantize the informativeness. Currently, several evaluation metrics (Zheng et al., 2024; Ke et al., 2023) based on LLMs have been proposed devoted to aligning model evaluation with human evaluation. Some works apply LLMs to the evaluation of long-form summarization, for example, Luo et al. (2023) uses ChatGPT to evaluate factual inconsistency for text summarization. More Recently, (Jung et al., 2024) calculates the saliency of a summary using MSM(Masked Language Model) to find how a candidate summary can be recovered given a reference summary. However, we claim that making use of the inherent ability of LLMs only is not enough to measure the ability of models to summarize for the lack of evaluation of the degree of information contained. Moreover, leveraging pre-trained probability distribution is not good enough because of its unexplainability. We hope to find a reasonable and explainable method for evaluation.
3 Preliminaries
In this section, we provide definitions for elements used in our evaluation pipeline.
Our work begins with extracting atomic facts from documents. Specifically, we denote reference summary as $s$ and candidate summary as $\hat{s}$. The extracted atomic fact lists are $S$ and $\hat{S}$ with $s_{i}$ and $\hat{s_{i}}$ being single atomic facts in the list. After atomic fact chain generation, we obtain lists containing logic chains $C$ and $\hat{C}$ for reference and candidate, while the elements of them are defined as $C_{i}$ and $\hat{C_{i}}$ and atomic fact in logic chains are defined as $C_{i_{j}}$ and ${\hat{C_{i_{j}}}}$. Finally, we conduct Document-level Relation Extraction(DocRE) to complement the original atomic fact list with hierarchical information. The final lists are $DocS$ and $Doc\hat{S}$ separately.
4 Methods
In this section, we introduce our proposed pipeline to evaluate informativeness in long-form summarization by extraction and validation. Given $s$ and $\hat{s}$, our automatic evaluation metric operates by 1) obtaining all atomic facts from $s$ and $\hat{s}$; 2) reforming atomic facts to form logic chains; 3) extracting all document-level relations and filter the duplicated one; 4) using LLM to validate whether it is true or false.
Figure 3: Examples show the generation of atomic facts. It uses demonstration to help LLMs understand the task.
4.1 Atomic Fact Generation
Atomic facts are defined as concise sentences that convey a single piece of information. In any given sentence within a model prediction, there may be multiple pieces of information that require individual verification. Consequently, evaluating the truthfulness of sentences on a one-by-one basis renders the true-or-false outcomes inadequate for assessing informativeness. Previous studies find that splitting candidate prediction into atomic facts (Min et al., 2023) improves the evaluation by a large margin and enables a more fine-grained evaluation. Following their work, we use ChatGPT with same prompt to generate atomic facts from both $S$ and $\hat{S}$ as $\left[s_{1},s_{2},s_{3}...\right]$ and $\left[\hat{s_{1}},\hat{s_{2}},\hat{s_{3}}...\right]$ where $s_{i}$ is atomic fact from reference summarization and $\hat{s_{i}}$ is atomic fact from candidate summarization.
4.2 Atomic Fact Chain Generation
Although claimed atomic, an examination of the “atomic facts” presented reveals that they often encompass more than one atomic fact. As illustrated in Figure 3, certain information cannot be independently extracted due to its reliance on contextual semantics. Consequently, in cases of cascaded atomic facts, where subsequent facts incorporate or entirely encompass information from preceding facts, if any of the preceding facts are incorrect, all subsequent facts are deemed incorrect, even if the newly added information is accurate.
To clarify, consider the examples: He played for Berekum Arsenal” and He played for Berekum Arsenal before moving to the United States.” These are two cascaded facts, with the latter not being atomic. However, the latter sentence cannot be decomposed into atomic facts because the adverbial clause depends on the entirety of the sentence for its meaning. Consequently, if the initial statement is incorrect, any subsequent statements, including those with additional correct information, are also considered incorrect.
We propose using LLMs, e.g. Mixtral-7B-instruct in our work to find logical relationships between atomic facts and treating atomic facts as a list of logic chains. The reformed lists are represented as $[C_{1},C_{2},C_{3}...]$ and $[\hat{C_{1}},\hat{C_{2}},\hat{C_{3}}...]$.
4.3 Document-level Relation Extraction
Extracting atomic facts from text primarily captures sentence-level relations. However, it is crucial to also consider document-level relations, which are challenging to discern solely through atomic fact generation. These relations are integral to understanding the contextual logic of a document. We complement Document-level relation extractions(DocRE) using the state-of-the-art model GPT-4 with in-context learning for both $s$ and $\hat{s}$ and the results are $[(h_{1},r_{1},t_{1}),(h_{2},r_{2},t_{2})...]$ and $[(\hat{h_{1}},\hat{r_{1}},\hat{t_{1}}),(\hat{h_{2}},\hat{t_{2}},\hat{r_{2}})...]$ where $(h_{i},r_{i},t_{i})$ and $(\hat{h_{i}},\hat{r_{i}},\hat{t_{i}})$ are head entity, relation and tail entity correspondingly. We then paraphrase the triples to make them natural language for post-processing.
To ensure that no sentence-level relations are included and contaminate work later, we use BERTScore along with TF-IDF to filter out the similar ones existing both in the atomic fact list and document-level relation list.
4.4 LLM for Validation
For each atomic fact $s_{i}$ in the candidate atomic fact list, we retrieve $k$ most relevant atomic facts from the reference summarization atomic fact list where k is a hyper-parameter which is a trade-off between accuracy and efficiency. For each atomic fact $\hat{C_{i_{j}}}$ in the corresponding logic chain $\hat{C_{i}}$, we check out one by one according to $s_{i}$ from $k$ most relevant atomic facts in $DocS$ represented as $f(\hat{C_{i_{j}}},s_{i})$. Specifically, we use Mixtral-7B-Instruct as our validation model. By providing the validation model with previous states and prompting it not to pay attention to previous information, we ensure that one piece of information is treated at a time.
We calculate our Precision and Recall as follows:
$\text{Precision}=\frac{1}{|Doc\hat{S}|}\sum_{i=0}^{n}\mathds{1}\left(f(\hat{C_{i_{j}}},s_{i})\right)$ (1)
$\text{Recall}=\frac{1}{|DocS|}\sum_{i=0}^{n}\mathds{1}\left(f^{-1}(s_{i},\hat{C_{i_{j}}})\right)$ (2)
where $n$ is the number of atomic facts in $Doc\hat{S}$. We define EVA-Score as the harmonic mean of Precision and Recall.
5 Experiment
5.1 Dataset
To test the applicability of EVA-Score, we collect long-form summarization with different context lengths. Specifically, we adopt CNN Dailymail(See et al., 2017), PubMed, GovReport(Cohan et al., 2018), arXiv, BookSum(Kryściński et al., 2022). In particular, we use the chapter-level summarization version of BookSum instead of the book level for less difficulty. For all these datasets, we use validation datasets for evaluation.
Dataset CNN Dailymail Pubmed arXiv Gov Report BookSum
Average Text Length 781 3111 5894 7797 14337
Average Summary Length 56 216 172 573 570
Table 1: The Average Text Length and the Average Summary Length in our experimented dataset.
5.2 Model
To perform a fair comparison, we experiment on a wide range of Generative Language Models of different numbers of parameters, and from different architectures. For open-source models, we test EVA-Score on Mistral-7B(Jiang et al., 2023) which is a MoE model, variants of Llama-2-chat(Touvron et al., 2023), with parameters 7B and 13B, variants of Vicuna(Zheng et al., 2023) with parameters 7B and 13B and Gemma-7B-it(Team et al., 2024). For close-source models, we experiment on ChatGPT and GPT-4(0125-preview version).
5.3 Baseline
We compare the results of EVA-Score with traditional evaluation metrics like ROUGE and BERTScore(Zhang et al., 2020) which are based on text overlap and word embeddings. Recently, many works have leveraged LLMs to obtain the informativeness of a given text corresponding to a reference. We include Vicuna-13B and Mixtral-7B-Instruct as our open-source LLM evaluators and ChatGPT and GPT-4(0125-preview version) as our close-source LLM evaluators. For all the LLMs, we instruct them using the prompt from Table 12, imitating the human annotation process.
5.4 Human Annotation
We are committed to developing an evaluation metric that resonates with human judgment. To achieve this, we randomly select 30 records from the dataset and assess the informativeness of each record by ourselves. Initially, we identify sentence-level and document-level atomic facts, denoted as $S$ and $\hat{S}$, from $s$ and $\hat{s}$, respectively. They then determine the overlap between $S$ and $\hat{S}$. Redundant information, which appears multiple times, is recorded only once. Subsequently, we calculate Precision and Recall for these overlaps and derive the F1 Score as their harmonic mean. Furthermore, we ensure that each record is evaluated by at least two annotators to maintain assessment fairness.
5.5 Main Results
Level Text-Level System-Level
Metric $\sigma$ $\gamma$ $\tau$ $\sigma$ $\gamma$ $\tau$
Traditional Evaluation Metric
ROUGE-1 0.305 0.160 0.114 0.455 0.800 0.667
ROUGE-2 0.413 0.203 0.127 0.014 0.399 0.333
BERT-Score 0.534 0.533 0.413 0.623 0.600 0.333
Model-based Evaluation Metric
Mixtral-7B-Instruct 0.097 0.052 0.030 -0.653 -0.400 -0.333
Vicuna-13B 0.065 -0.040 -0.029 0.367 0 0
ChatGPT 0.138 0.210 0.162 0.430 0.600 0.333
GPT-4 0.308 0.329 0.229 0.505 0.600 0.333
EVA-Score (Ours) 0.710 0.673 0.503 0.975 1.000 1.000
Table 2: Text-level and system-level Pearson($\sigma$), Spearman($\gamma$), and Kendall($\tau$) correlations with different evaluation metrics. The highest correlation among the methods is bold
In this section, we report the correlation between each evaluation metric and human annotations. We adopt text-level and system-level Pearson($\sigma$), Spearman($\gamma$), and Kendall($\tau$) correlation. Text-level correlation is calculated by computing the correlation coefficients between human evaluations and automatic evaluation metrics and then averaging these coefficients for all generated texts per instruction. System-level correlation, on the other hand, is determined by computing the correlation coefficients between human evaluations and automatic evaluation metrics, but this time for the average score of each dataset.
From Table 2, we observe that our EVA-Score outperforms all other model-based and traditional evaluation metrics in achieving the highest correlation with human judgments across all indicators at both levels. At the System Level, EVA-Score achieves perfect Spearman and Kendall correlations of 1.00, demonstrating complete alignment with human assessments. In contrast, Traditional Evaluation Methods, while somewhat correlated with human evaluations, generally lack strong correlations, highlighting their limitations.
Surprisingly, we find that the information scores assigned by Vicuna-13B show little correlation with humans, and At the System Level, the scores from Mistral-7B show even negative correlation with humans. We attribute this to the situation that the informativeness is obtained by quantitative calculation using EVA-Score and Human Annotation. The instruction may be too complex for them to deal with within a single interaction which requires them first to extract atomic facts and document-level relations and verify all of them one by one. However, for models like ChatGPT and GPT-4, while their correlation with human judgments at the Text Level remains minimal, they exhibit promising outcomes at the System Level using only In-Context Learning.
We also report our results on several long-form summarization datasets, including CNN Dailymail, PubMed, GovReport, arXiv, and BookSum, using three different evaluation metrics, i.e. EVA-Score, ROUGE-1, and BERTScore. Since this work focuses on the evaluation of long-form summarization, we don’t pursue feeding long context into LLMs without information loss. Instead, we believe that useful information mostly appears at the beginning or the end of a text, we just simply truncate the context from both the beginning and end to meet the models’ max input sequence length.
Dataset CNN Dailymail Pubmed arXiv Gov Report BookSum
Metric EVA R-1 BS EVA R-1 BS EVA R-1 BS EVA R-1 BS EVA R-1 BS
Vicuna-7B 0.501 0.441 0.873 0.327 0.398 0.832 0.327 0.344 0.838 0.239 0.308 0.846 0.145 0.212 0.825
Vicuna-13B 0.462 0.432 0.872 0.431 0.388 0.837 0.346 0.348 0.832 0.154 0.298 0.849 0.144 0.188 0.827
Mixtral-7B-Instruct 0.499 0.422 0.872 0.447 0.447 0.847 0.444 0.448 0.845 0.313 0.399 0.854 0.259 0.377 0.839
Llama-2-7B-chat 0.399 0.386 0.868 0.358 0.421 0.845 0.411 0.435 0.851 0.193 0.198 0.798 0.186 0.268 0.826
Llama-2-13B-chat 0.453 0.423 0.868 0.416 0.405 0.841 0.444 0.332 0.830p 0.296 0.378 0.857 0.154 0.280 0.827
Gemma-7B-it 0.431 0.383 0.858 0.492 0.417 0.842 0.385 0.359 0.831 0.250 0.291 0.843 0.154 0.256 0.818
ChatGPT 0.472 0.373 0.864 0.444 0.359 0.838 0.443 0.368 0.838 0.302 0.261 0.854 0.182 0.202 0.826
GPT-4 0.466 0.376 0.863 0.454 0.407 0.838 0.453 0.361 0.835 0.325 0.332 0.854 0.257 0.261 0.824
Table 3: The results of Several LLMs on different long-form summarization datasets. We denote EVA-Score as EVA, R-1 as ROUGE-1, BS as BERTScore. All the values are normalized between 0 and 1. We indicate the highest score of EVA-Score using bold.
Referring to Table 3, we observe that as the context lengthens, all tested models experience varying degrees of performance decline. Particularly notable is the performance at the part-of-book level, i.e. chapter summarization in BookSum, where the results are markedly poor compared to the ground truth. This issue, however, can be mitigated by employing methods that feed the entire context into the model, rather than truncating it as we have done in our initial experiments. Additionally, compared to other evaluation metrics, the EVA-Score demonstrates a more pronounced and consistent decline in performance as context length increases, with fewer fluctuations. This pattern aligns well with human cognitive processes, indicating that as information load increases, the challenge of maintaining high summarization quality also grows. This consistent pattern of decline not only reflects realistic expectations of model performance but also underscores the effectiveness of the EVA-Score in capturing these dynamics, further validating its utility from another perspective.
Moreover, models using ROPE(Su et al., 2023) or other methods, enabling longer input sequences show greater performance in long-form summarization when the context length becomes larger, mainly due to their ability to comprehend a long context or search in a long context. However, when the context is not so long, they show comparable or inferior performance with assistant models with normal sequence length, e.g. 4096 or 8192.
Under the evaluation framework of the EVA-Score, GPT-4 demonstrates robust capabilities, particularly with longer input sequences, compared to other models. In the CNN Dailymail dataset, it trails behind models such as Vicuna-7B and ChatGPT. This is primarily due to GPT-4’s tendency to generate more extensive summaries than its counterparts in settings with relatively short contexts, thereby increasing its count of atomic facts. While this might allow it to cover more points than other models, it often results in lower precision due to the inclusion of information that extends beyond the reference summary. Nevertheless, in other datasets such as arXiv and Gov Report, GPT-4 outperforms all competitors. It also secures a second-place ranking in Pubmed and BookSum, affirming its effectiveness across diverse text summarization tasks.
To be mentioned, Mixtral-7B-Instruct, compared with other models with the same parameter or even larger, shows great potential to fulfill long-form summarization and even beats GPT-4 by a small margin in chapter-level BookSum, which meets the findings of Chang et al. (2024).
5.6 Analysis of Summarization Length
Dataset CNN Dailymail Pubmed arXiv Gov Report Booksum
Indicator Avg Length Avg Length Avg Length Avg Length Avg Length
Vicuna-7B 108 214 133 258 124
Vicuna-13B 106 146 205 188 99
Mistral-7B-Instruct 109 207 212 353 251
Llama-2-7B-chat-hf 118 143 108 140 185
Llama-2-13B-chat-hf 124 184 221 236 179
Gemma-7B-it 98 139 140 177 151
ChatGPT 102 118 124 137 107
GPT-4 153 227 272 242 212
Table 4: The average summarization length(denoted by Avg Length by different model. The longest is marked underline
To elucidate why GPT-4 underperforms on the EVA-Score with a shorter context, we conduct an analysis focused on the summarization lengths produced by different models, as detailed in Table 4. Models pre-trained with longer input sequences tend to generate more extensive responses. Specifically, GPT-4 produces summaries of 153 tokens in CNN Dailymail, 227 tokens in Pubmed, and 272 tokens in arXiv and Mixtral-7B-Instruct outputs 353 tokens in Gov Report and 251 tokens in BookSum whose maximum input sequences length are 128000 and 32768 respectively.
Upon manual examination, we observe that shorter contexts typically result in shorter ground truth. Consequently, if a model generates a lengthy prediction, this verbosity is penalized with a reduced EVA-Score due to diminished precision. As the context lengthens, references become richer in information. However, these models tend to maintain a similar response length, lacking significant information addition, thereby failing to capture enough relevant points, especially in longer predictions. This discrepancy explains GPT-4’s shortfall in CNN Dailymail and the superior performance of Mistral-7B-Instruct, despite having a comparable number of parameters.
Furthermore, GPT-4, which is refined through Reinforcement Learning from Human Feedback (RLHF), typically caps its responses to summarizations under 300 tokens, even in contexts of substantial length. This behavior is a contributing factor to GPT-4’s underachievement in BookSum, where more extensive and detailed responses are crucial for capturing the full scope of the content.
5.7 Ablation Study
We report the effect of Atomic Fact Chain Generation and Document-level Relation Extraction on the number of atomic facts to be verified.
Settings AFG only AFCG AFCG + DocRE
CNN Dailymail 25 19 24
Pubmed 38 25 30
arXiv 33 22 26
Gov Report 44 26 30
BookSum 36 27 31
Table 5: Number of facts(or fact chains) to be evaluated under three settings, averaging 9 models we used for experiments, i.e. Atomic Fact Generation(AFG), Atomic Fact Chain Generation(AFCG) and Atomic Fact Chain Generation Complemented with DocRE(AFCG + DocRE).
From Table 5, datasets with longer context and richer information tend to result in more facts, and by using Atomic Fact Chain Generation, we reduce the number of cascaded facts a lot. By doing so, minor mistakes in one piece of information will not affect others. Also, DocRE complements AFG with document-level relations, which is helpful for the evaluation of the hierarchical information in the summary.
5.8 Meta Evaluation of Human Annotation
By observing the results of human annotations, we find that ground truth prefers to summarize more elegantly, using core arguments from the original context in addition to examples supported with other a little bit trivial arguments. However, LLMs try to list all of the arguments, ignoring the examples or data in the context, leading to differences with ground truth. However, there does exist some verbosity in the ground truth, indicating the natural drawback of the summarization dataset.
5.9 Analysis of Different LLM Verifier
Since our LLM Verifier should consider previous states in a logic chain and verify whether the current atomic fact is true or not, it is a rather hard problem compared with simply verifying. Intuitively, with more capable models, we will get a more human-aligned evaluation metric. We report the results using ChatGPT for a simple proof.
Metric $\sigma$ $\gamma$ $\tau$
EVA-Score-Mistral-7B-Instruct 0.710 0.673 0.503
EVA-Score-ChatGPT 0.737 0.672 0.543
Table 6: Human correlation using ChatGPT as EVA-Score LLM verifier compared with EVA-Score with Mistral-7B-Instruct. The larger one is denoted is bold.
Leveraging ChatGPT results in more human correlation in Pearson and Kendall correlation and comparable Spearman correlation, which is consistent with our recognition. However, using ChatGPT results in much more costs than open-source LLMs because of the great number of validations.
5.10 Case Study
To provide a more intuitive and straightforward insight into the pipeline of EVA-Score, we provide a detailed case in Appendix B. From the case, we can see that EVA-Score provides not only a quantitative and objective information score but also provides a human-like explanation.
6 Conclusion and Future Work
In this work, we propose a new evaluation metric called EVA-Score to evaluate informativeness in long-form summarization. It extracts atomic facts which are then reformed as logic chains and document-level relations from reference and candidate and uses LLMs for validation one by one. Experiment results show that compared with other traditional metrics or LLM-based metrics, our work shows the best human correlation as well as provides an objective explanation. Moreover, we re-evaluate long-form summarization based on EVA-Score, indicating a great challenge in book-level summarization.
We believe that EVA-Score is meant to calculate information overlap between reference and candidate, thus it can be applied to broader scenarios rather than long-form summarization only. In future work, we suggest: 1) Leveraging EVA-Score to evaluate more reference-based and information-rich tasks and 2) Utilizing EVA-Score for the generation of more informative texts.
Limitations
EVA-Score can be time-consuming to run
EVA-Score includes several pre-processing methods. We should first extract atomic facts and document-level relations two summaries and use LLMs for paraphrasing to make the relations more natural. The pre-processing takes around two minutes per record. Then for each atomic fact in the candidate, we select $k$ most relevant from the reference and use LLMs for validation. Since the number of atomic facts takes quite a lot of time. We report the time consumed correlated to the sum of the length of the candidate and reference.
Length Interval/token 0-300 300-600 600-900 900-1200
Time Consuming/s 195 402 578 629
Table 7: Time consumed regarding to different summary length
From Table 7, for a summary of relatively short length, it takes about three minutes to finish the evaluation, while for a longer summary, it may take about ten minutes which is a long time compared to traditional methods.
EVA-Score only considers Informativeness
Some existing evaluation metrics consider many other factors, like faithfulness, and consistency. EVA-Score may be combined with other evaluation metrics to provide a more comprehensive evaluation of long-form summarization.
EVA-Score Leverages GPT-4 for Doc-RE
EVA-Score implements Document-level Relation Extraction for the lack of hierarchical information. However, Doc-RE by GPT-4 maybe incomplete because GPT-4 may fail to capture the core meaning of the document level due to the difficulty of such a task. Leveraging more capable models that are fine-tuned on Doc-RE may be more helpful.
Acknowledgments
This work is supported by the National Science and Technology Major Project(2023ZD0121403). We extend our gratitude to the anonymous reviewers for their insightful feedback.
Appendix A Instructions for EVA-Score
A.1 Instruction for Atomic Fact Generation
We use the following prompt to generate atomic facts.
Instruction: Please breakdown the following sentence into independent facts:
Domonstrations
Example 1
He made his acting debut in the film The Moon is the Sun’s Dream (1992), and continued to appear in small and supporting roles throughout the 1990s.
- He made his acting debut in the film.
- He made his acting debut in The Moon is the Sun’s Dream.
- The Moon is the Sun’s Dream is a film.
- The Moon is the Sun’s Dream was released in 1992.
- After his acting debut, he appeared in small and supporting roles.
- After his acting debut, he appeared in small and supporting roles throughout the 1990s.
Example 2 In 1963, Collins became one of the third group of astronauts selected by NASA and he served as the back-up Command Module Pilot for the Gemini 7 mission.
- Collins became an astronaut.
- Collins became one of the third group of astronauts.
- Collins became one of the third group of astronauts selected.
- Collins became one of the third group of astronauts selected by NASA.
- Collins became one of the third group of astronauts selected by NASA in 1963.
- He served as the Command Module Pilot.
- He served as the back-up Command Module Pilot.
- He served as the Command Module Pilot for the Gemini 7 mission.
Table 8: The instruction for atomic fact generation
A.2 Instructon for Atomic Fact Chain Generation
As discussed before, we use the following prompt to decide whether one fact is entailed by others.
Instruction: Given a hypothesis and a reference, determine if the hypothesis is entailed by the reference. Entailed means that if we use the information from reference, we can infer the hypothesis, e.g. the hypothesis is in a logical relationship with the reference. If the hypothesis is entailed by the reference, answer True. If the hypothesis contradicts the reference or the hypothesis is neural,answer False. If the answer is ambiguous, answer False. The output should be True and False only. You should not explain why you made this choice. Now let’s begin the question:
Domonstrations
Example 1 Hypothesis: The cat is on the mat. Reference: The cat is sleeping on the mat. Answer: True Example 2 Hypothesis: The cat is on the mat. Reference: The dog and the rat are on the mat. Answer: False
Table 9: The instruction for atomic fact chain generation
A.3 Instruction for Document-level Relation Extraction
We use the following prompt with GPT-4 to conduct Document-level Relation Extraction.
Instruction: Given a text and entities in the given text, your task is to extract document-level relation triples from the text. A relation triple is a 3-tuple (subject, relation, object) where subject and object are named entities and relation is the relation between them. For example, given the text "Bill Gates is the founder of Microsoft", the extracted relation triple would be (Bill Gates, founder, Microsoft). In document-level RE, subject and object entities in a given triple might be dispersed across distinct sentences, and certain entities may have aliases in the form of distinct entity mentions. Therefore, you should pay attention to relations across sentences and entity aliases rather than sentence-level relations. i.e. You should not respond with sentence-level relations, but document-level relations. The answer format is listed below, you should obey it for any situation. Any additional information should not be added. Neither head entity nor tail entity should be empty strings. The relation triples should be in the following format: (’head entity’, ’relation’, ’tail entity’) (’head entity’, ’relation’, ’tail entity’) (’head entity’, ’relation’, ’tail entity’)
Demonstrations
Text
[…]
Entities
[…]
(’Bucharest’, ’country’, ’Romania’)
(’Bucharest’, ’located in the administrative territorial entity’, ’Romania’)
(’Romanian Air Force’, ’country’, ’Romania’)
(’Aurel Vlaicu Airport’, ’country’, ’Romania’)
(’Aurel Vlaicu Airport’, ’located in the administrative territorial entity’, ’Bucharest’)
(’Henri Coandă International Airport’, ’country’, ’Romania’)
(’Urziceni’, ’country’, ’Romania’)
(’Radu Berceanu’, ’country of citizenship’, ’Romania’)
[…]
(’Alexeni’, ’country’, ’Romania’)
Table 10: The instruction for Document-level Relation Extraction
A.4 Instruction for LLM validation
We use the prompt below to instruct LLMs focus on one piece of information each time.
Instruction: Given a hypothesis and a reference, determine if the hypothesis is entailed by the reference. Entailed means that if we use the information from reference, we can infer the hypothesis, e.g. the hypothesis is in a logical relationship with the reference. The hypothesis may be part of a logical chain, so we may provide you with the correctness of previous states. If we provide you with the correctness of previous states, you should not focus on the information that is redundant, that is, the information that exists both in previous facts and the current fact. Whether they are correct or not, you should not care them. You should only concentrate on new information. If the hypothesis is entailed by the reference, answer True. If the hypothesis contradicts the reference or the hypothesis is neural,answer False. If the answer is ambiguous, answer False. The output should be True and False only. You should not explain why you make this choice.
Table 11: The instruction for LLM validation
A.5 Instruction for LLM-based Evaluation of Informativenss
We use the following prompt in Table 12 to instruct LLMs to generate the informative score.
Instruction:
Given a Reference Answer and a Candidate Answer, your task is to first extract all information important to the Reference Answer and the Candidate Answer, and then calculate the Precision, Recall, and F1-score of the Candidate Answer Information pool concerning the Reference Answer Information pool. Information is defined as words, phrases or sentences that are important to the Reference Answer and the Candidate Answer. The subjective, objective, adverbial and attribute may all be included in the information. The information pool is the collection of all information extracted from the Reference Answer and the Candidate Answer. You can use the following function to calculate the Precision, Recall, and F1-score of the Candidate Answer Information pool with respect to the Reference Answer Information pool. def calculate_metrics(gold_key_information_pool, model_key_information_pool, matching_pool):
    precision = len(matching_pool) / len(gold_key_information_pool)
    recall = len(matching_pool) / len(model_key_information_pool)
    f1_score = 2 * precision * recall / (precision + recall)
    return f1_score
Finally, only return with a floating point number rounded to 3 decimal places.
Demonstrations
Reference Summary
background: coronary computed tomography angiography ( ccta ) is a frequently performed examination for coronary artery disease . when performed with retrospective gating , there is an opportunity to derive functional parameters of left ventricle utilizing automated software . complementary information , if validated with established standards , will enhance the total value of study.objective:study evaluates the usefulness of fully automated software for the assessment of left ventricular ejection fraction ( lvef ) using 64-slice ccta data and to correlate ct results with echocardiography ( echo ) . role of ct derived lv function is reviewed in the light of emerging technologies and recent developments in multidetector ct ( mdct).materials and methods : a total of 113 patients referred for mdct ccta for evaluation of coronary artery disease . all patients were scanned on 64 slice ge - helical ct scanner and had an echo done within 1 week of the ct scan . retrospectively electrocardiogram ( ecg)-correlated image reconstruction was performed with the reconstruction at 10% r - r interval increment . axial image sets were analyzed with advanced workstation using a program - auto ejection fraction , circulation : ge medical solutions.results:the mean lvef calculated by clinical echo was 58.6 4.5% and by fully automated software based on cta data was 58.9 5.4% . the pearson ’s regression analysis showed a large correlation , with a correlation coefficient of 0.503 ( p < 0.001 ) . bland - altman analysis showed a trend towards mdct resulting in slightly higher values for lvef when compared with echo.conclusion:the fully automated software is simple , reliable , and user - friendly , and can provide rapid assessment of lv functional parameters with good reproducibility . despite of good correlation , fewer patients are likely to benefit , in future , from this function due to smaller number of patients undergoing ccta with retrospective gating Candidate Summary
The text summarizes the results of a study on the use of multidetector computed tomography (MDCT) for assessing left ventricular (LV) function in patients undergoing coronary angiography. The study found that MDCT can provide useful functional information in addition to coronary angiography, and that fully automated analysis software can be used to measure LV ejection fraction (EF) with good reproducibility. However, the study also noted that the radiation dose involved in MDCT is a limitation, and that additional clinically indicated coronary CT examination with a lowest possible radiation dose is invaluable in patients known or suspected of ischemic heart disease. The study also found that validating consistency of results with magnetic resonance imaging (MRI) will further lend ‘ to the use of MDCT derived results. Thought Gold key information pool = [coronary computed tomography angiography, examination, coronary artery disease, retrospective gating,derive functional parameters,utilizing automated software, validated with established standards,enhance, total value, evaluates the usefulness, fully automated software, left ventricular ejection fraction, 64-slice ccta data, ct results, echocardiography ( echo ), reviewed, emerging technologies,multidetector ct ( mdct), 113 patients, evaluation, scanned,helical ct scanner, 1 week of,electrocardiogram, performed with the reconstruction, 10%r - r interval increment, axial image,analyzed with advanced workstation,auto ejection fraction,the mean, 58.6 4.5%, cta data was 58.9 5.4%, pearson ’s regression analysis,large correlation,coefficient of 0.503 ( p < 0.001 ), altman analysis,slightly higher values for lvef, compared, friendly,rapid assessment, good reproducibility, fewer patients, reliable, lvef] Model Key information pool = [multidetector computed tomography (MDCT), assessing left ventricular (LV), patients, coronary angiography, provide,useful functional information, automated analysis software, Measure, ejection fraction (EF), good reproducibility, radiation dose, limitation, clinically indicated, coronary CT examination, lowest possible, radiation dose, invaluable, patients known,suspected, ischemic heart disease, validating consistency, magnetic resonance imaging (MRI), further lend support, derived] Matching pool: [multidetector computed tomography (MDCT), assessing left ventricular (LV) match assessing left ventricular (LV), patients match patients, coronary angiography match coronary angiography, coronary angiography match coronary angiography, ejection fraction match lvef, good reproducibility match good reproducibility, slightly higher values match invaluable, derive functional parameters match derived]
F1 = calculate_metrics(gold_key_information_pool, model_key_information_pool, matching_pool)
print(F1) Output:
0.228
Table 12: The instruction for Document-level Relation Extraction
Appendix B Case Study
To provide a more intuitive and straightforward insight into the pipeline of EVA-Score, we provide a detailed case. From the case, we can see that EVA-Score provides not only a quantitative and objective information score but also provides a human-like explanation.
Reference Summary:
The mpp decreased the number of dopaminergic neurons by 40% , and increased the release of lactate dehydrogenase ( ldh ) into the culture medium . The tq significantly rescued dopaminergic neurons and decreased the release of ldh at the concentrations of 0.1 and 1 m
Atomic Fact Chains
[- The mpp decreased dopaminergic neurons, - The mpp decreased the number of dopaminergic neurons, - The mpp decreased the number of dopaminergic neurons by 40%]
[- The mpp increased the release of lactate dehydrogenase ( ldh ), - The mpp increased the release of lactate dehydrogenase ( ldh ) into the culture medium]
[- The tq significantly rescued dopaminergic neurons]
[- The tq decreased the release of ldh, - The tq decreased the release of ldh at the concentrations, - The tq decreased the release of ldh at the concentrations of 0.1 and 1 m]
Document-level Relations:
tq causes Thymoquinone
mpp leads to 1-methyl-4-phenylpyridinium
Candidate Summary:
The MPP significantly decreased the survival of dopaminergic neurons and increased the release of lysosomal hydrolase (LDH) into the culture medium. The neurotoxicity of MPP was attributed to its selective uptake by dopaminergic neurons through the dopamine transporter.
Atomic Fact Chains
[- The MPP significantly decreased dopaminergic neurons, -The MPP significantly decreased the survival of dopaminergic neurons ]
[- The MPP increased the release of lysosomal hydrolase (LDH), - The MPP increased the release of lysosomal hydrolase (LDH) into the culture medium]
[- The neurotoxicity was attributed to dopaminergic neurons, - The neurotoxicity of MPP was attributed to dopaminergic neurons, - The neurotoxicity of MPP was attributed to its selective uptake by dopaminergic neurons, - The neurotoxicity of MPP was attributed to dopaminergic neurons through the dopamine transporter, - The neurotoxicity of MPP was attributed to its selective uptake by dopaminergic neurons through the dopamine transporter.]
Document-level Relations:
mpp causes 1-methyl-4-phenylpyridinium
neurotoxicity infects neuroprotective effect
Matching Process
- “The mpp decreased dopaminergic neurons” match “The MPP significantly decreased dopaminergic neurons”
- “The mpp increased the release of lactate dehydrogenase ( ldh )” match “The mpp increased the release of lactate dehydrogenase ( ldh )”
- “The mpp increased the release of lactate dehydrogenase ( ldh ) into the culture medium” match “The MPP increased the release of lysosomal hydrolase (LDH) into the culture medium”
- “The mpp decreased dopaminergic neurons and increased the release of lactate dehydrogenase ( ldh )” match ‘The MPP significantly decreased dopaminergic neurons and increased the release of lysosomal hydrolase (LDH)”
- “The mpp decreased dopaminergic neurons and increased the release of lactate dehydrogenase ( ldh ) into the culture medium” match “The MPP significantly decreased dopaminergic neurons and increased the release of lysosomal hydrolase (LDH) into the culture medium”
- “mpp causes 1-methyl-4-phenylpyridinium” match “mpp leads to 1-methyl-4-phenylpyridinium”
Table 13: Case Study For EVA-Score from Pubmed Dataset.

### v3

@@
-EVA-Score: Evaluation of Long-form Summarization on Informativeness through Extraction and Validation
+EVA-Score: Evaluating Abstractive Long-form Summarization on Informativeness through Extraction and Validation
 Abstract
-Summarization is a fundamental task in natural language processing (NLP) and since large language models (LLMs), such as GPT-4 and Claude, come out, increasing attention has been paid to long-form summarization whose input sequences are much longer, indicating more information contained.
-The current evaluation metrics either use similarity-based metrics like ROUGE and BERTScore which rely on similarity and fail to consider informativeness or LLM-based metrics, lacking quantitative analysis of information richness and are rather subjective.
-In this paper, we propose a new evaluation metric called EVA-Score using Atomic Fact Chain Generation and Document-level Relation Extraction together to automatically calculate the informativeness and give a definite number as an information score.
-Experiment results show that our metric shows a state-of-the-art correlation with humans.
-We also re-evaluate the performance of LLMs on long-form summarization comprehensively from the information aspect, forecasting future ways to use LLMs for long-form summarization.
-EVA-Score: Evaluation of Long-form Summarization on Informativeness through Extraction and Validation
-Yuchen Fan1thanks:  The work was done when he worked as an intern at Tsinghua University, China., Xin Zhong1, Chengsi Wang1, Gaochen Wu2, Bowen Zhou2thanks:  Corresponding authors 1 Beijing University of Posts and Telecommunications 2 Department of Electronic Engineering, Tsinghua University yuchenfan48@gmail.com zhoubowen@tsinghua.edu.cn
+Since LLMs emerged, more attention has been paid to abstractive long-form summarization, where longer input sequences indicate more information contained.
+Nevertheless, the automatic evaluation of such summaries remains underexplored.
+Current evaluation metrics either use traditional metrics like ROUGE and BERTScore, which rely on surface-level similarity and fail to consider informativeness, or simple LLM-based metrics, which are not robust and easily overwhelmed by the long contexts.
+In this paper, we propose a new evaluation metric called EVA-Score to extract all information from the given summaries, identify overlapped information based on reference, and calculate the information score.
+We test EVA-Score on several datasets and the experimental results reveal that EVA-Score shows the highest correlation with humans.
+We also re-evaluate the performance of LLMs on long-form summarization from the information perspective.
+The results indicate that responses of LLMs still have a gap with the human-written answers.
+Moreover, we provide a detailed analysis of the effectiveness of EVA-Score, forecasting future ways to automatically evaluate abstractive long-form summarization.
+EVA-Score: Evaluating Abstractive Long-form Summarization on Informativeness through Extraction and Validation
+Yuchen Fan1,thanks:  The work was done when he worked as an intern at Tsinghua University, China., Xin Zhong1, Yazhe Wan1, Chengsi Wang1, Haonan Cheng1, Gaochen Wu2, Ning Ding2, Bowen Zhou2,thanks:  Corresponding authors 1 Beijing University of Posts and Telecommunications 2 Department of Electronic Engineering, Tsinghua University yuchenfan48@gmail.com zhoubowen@tsinghua.edu.cn
 1 Introduction
-Figure 1: Examples show the dilemma of ROUGE.
-The blue annotation is the atomic information contained in the text.
-Even though Candidate Summarization 1 is rewarded with a higher ROUGE Score, it contains less information than Candidate Summarization 2.
-It inspires us that ROUGE and BERTScore are not reasonable for evaluating summarization.
-Summarization has long been a fundamental task in NLP and since LLMs came out (OpenAI et al., 2024; Touvron et al., 2023), many research leverages LLMs for summarization (Zhang et al., 2023b; Wang et al., 2023b).
-More recently, works on summarization moved to Long-form scenarios where the input sequence length is extremely larger, resulting in more information contained in it.
-Traditional evaluation metrics may be not suitable for evaluation of long-form summarization due to the great informativeness, thus calling for new evaluation metrics.
-Past evaluation metrics for summarization include some static evaluation methods, like ROUGE-k and ROUGE-l (Lin, 2004) which calculate the k-gram overlap or the longest common substring between the reference and the candidate.
-With the development of pre-trained language models, BERTScore (Zhang et al., 2020) is widely used as a new evaluation metric based on pre-trained word embeddings.
-More recently, some works have begun to pay attention to evaluating long-form summarization (Kim et al., 2024; Chang et al., 2024) which uses prompts to define criteria to assign different scores for different aspects, like fluency, consistency, and coherence.
-However, traditional and LLM-based evaluation metrics face their dilemmas respectively.
-Figure 2: Pipeline of our evaluation metric.
+Figure 1: The example shows the dilemma of traditional evaluation metrics, including ROUGE and BERTScore.
+Candidate Summary 1 and Candidate Summary 2 are summaries to be evaluated.
+The information within the summary is denoted using blue.
+Even though Candidate Summary 1 is rewarded with a higher ROUGE and BERTScore, it contains less information and is given a lower EVA-Score which meets human preference.
+It inspires us that traditional metrics are not reasonable for evaluating long-form and knowledge-intensive summaries due to their insensitivity to information.
+Long-form summarization is challenging due to its extensive knowledge and lengthy context.
+It can be categorized into extractive (Zhang et al., 2023) and abstractive (Wang et al., 2023b), where the former retains exact content from the document, while the latter produces a paraphrased and coherent output.
+With advancements in LLMs (OpenAI et al., 2024; Touvron et al., 2023), significant progress has been made (Chang et al., 2024; Chhikara et al., 2024; Kim et al., 2024).
+However, the automatic evaluation of abstractive long-form summarization remains underexplored and is more challenging than evaluating extractive ones due to the implicit equivalence of information involved.
+Therefore, an information-sensitive metric is needed to effectively capture the depth and breadth of summaries.
+Past evaluation metrics for abstractive summarization include some static evaluation methods, like ROUGE-k and ROUGE-L (Lin, 2004).
+With the development of pre-trained language models, BERTScore (Zhang et al., 2020) is widely used as an evaluation metric based on word embeddings to measure the similarity between the reference and the candidate.
+Also, some works use LLMs to evaluate long-form summarization, defining criteria, and prompting LLMs to assign scores for different aspects like fluency and consistency.
+For clarity, we categorize metrics that utilize prompts and obtain responses from LLMs as LLM-based evaluation metrics (Kim et al., 2024; Chang et al., 2024), while other metrics like ROUGE and BERTScore are referred to as traditional metrics.
+Each category, however, has its limitations and challenges, especially when applied to complex, nuanced content in abstractive long-form summarization.
+Figure 2: Pipeline of EVA-Score.
 We begin with Atomic Fact Generation.
-For generated facts, we use LLMs to check whether they can be treated as a logic chain and use Document-level Relation Extraction as a complement.
-Finally, when analyzing the truthfulness of a given fact, we use retrieval to find several reference facts and provide previous states.
-For Similarity-based Metrics Since they are based on similarity, they cannot represent informativeness and are unsuitable in such an LLM era, where responses from LLMs are flexible and contain many synonyms and different syntax.
-For LLM-based Metrics They are insensitive to nuanced differences existing in summaries and their quantitative results are not reliable even if given detailed criteria for evaluation due to the black-box property.
-We propose a new evaluation metric whose result is an objective and quantitative analysis but can fully catch the information contained in the summary.
-We call this metric Extraction-and-Validation Score.
-We begin with extracting atomic facts from reference and candidate summaries.
-Different from FActScore (Min et al., 2023), we claim that some information cannot live without a context, so we treat them as part of logic chains.
-Afterward, we introduce Document-level Relation Extraction to remedy the loss of hierarchical relations.
-Then for each atomic fact to be verified, we use information retrieval to find the most relevant atomic facts as references and prompt LLMs to focus only on newly added information and decide whether it is entailed or not.
-Intuitively, LLMs for evaluation can be more accurate in single verification than in complex situations.
-We ensure that only one piece of information is evaluated at a time, simplifying the validation process.
-And by introducing Document-level Relations, we form a more comprehensive information set, maintaining the completeness of the evaluation.
-We conduct comprehensive experiments by comparing EVA-Score with current evaluation metrics, including traditional and LLM-based.
-The results show that EVA-Score achieves the best human alignment compared to SOTA past evaluation metric, BERTScore by 0.166, 0.120, and 0.09 in Pearson, Spearman, and Kendall correlation respectively at the text level.
-At the system level, it shows human-comparable evaluation, with 0.975 in Pearson correlation and 1.0 in both Spearman and Kendall correlation.
-We also re-evaluate long-form summarization based on EVA-Score and the experiment results show the informativeness of different LLMs on long-form summarization.
-The results show the great performance of GPT-4 when the context becomes longer, ranking first in arXiv and Gov Report and second in Pubmed and BookSum.
-It also shows the promising performance of Mixtral-7B-Instruct for long-form summarization.
-Our main contributions are as follows:
-  • We propose a new evaluation metric that can generate high-quality evaluation results of a given summary, based on atomic fact chain generation and document-level relation extraction, by providing a numerical score and steps to obtain this score.
-  • We conduct comprehensive experiments on several long-form summarization datasets, with the number of tokens of each record ranging from 1k to 14k.
-Experiment results show that our evaluation metric shows superiority in every correlation with humans.
-  • We re-evaluate long-form summarization concerning informativeness which is the first to evaluate from the information aspect as far as we know.
-In our experiments, the state-of-the-art model GPT-4 still performs better than any other models in long context settings but lags behind some assistant models in short context settings.
+For extracted facts, we use LLMs to check whether they can be treated as a logic chain and use Document-level Relation Extraction as a complement.
+“Doc Re” in the figure is actually “Document-level Relation” for abstraction.
+When analyzing the truthfulness of a given fact, we retrieve and find several reference facts and provide the evaluation results of previous facts within a logic chain for help.
+Thus, we ensure that one piece of information is analyzed once.
+Traditional metrics, such as ROUGE and BERTScore, are primarily based on surface-level similarity—either word-level or embedding-level similarity to the reference (Krishna et al., 2021).
+Therefore, they often fail to capture information-level similarity, which involves understanding the actual meaning or key ideas conveyed, making them less suitable in the era of LLMs, where responses are highly flexible and often use synonyms or different syntactic structures.
+We show an example in Figure 1, indicating the inherent drawback of ROUGE and BERTScore.
+On the other hand, LLM-based Metrics struggle with capturing nuanced information differences between candidates and references.
+Also, their evaluations can be unreliable, even with detailed evaluation criteria in such a complex scenario, where LLMs must first identify key information from the candidate and then assess it against the reference (Shen et al., 2023).
+To fill the gap, we propose a new evaluation metric called Extraction-and-Validation Score (EVA-Score), designed to provide an objective and quantitative analysis while capturing the full range of information in a summary.
+Our approach begins by extracting atomic facts from reference and candidate.
+Unlike FActScore (Min et al., 2023), we recognize that certain pieces of information are context-dependent, resulting in some extracted facts not being atomic but cascaded between them instead.
+To address this, we reformat these atomic facts into logic chains that preserve contextual integrity.
+Additionally, since atomic fact extraction operates on a sentence-level granularity, it often misses document-level relationships.
+For complement, we incorporate Document-Level Relation Extraction (Xue et al., 2024; Xu et al., 2022) to capture the hierarchical relations and provide a more complete understanding of the content.
+To verify the accuracy of each atomic fact, we identify the most relevant context from the reference and prompt LLMs to determine whether the candidate information is entailed by the reference.
+In our implementation, we ensure that only one piece of information is evaluated at a time, which simplifies the validation and enhances reliability.
+The overall evaluation pipeline is illustrated in Figure 2.
+To the best of our knowledge, we are the first to evaluate long-form summarization from the informativeness perspective.
+We conduct comprehensive experiments to compare EVA-Score with existing evaluation metrics, including both traditional and LLM-based metrics.
+The results reveal that EVA-Score achieves the highest alignment with human evaluations, significantly outperforming past metrics.
+At the system level, EVA-Score demonstrates a strong correlation with human judgments, achieving a Pearson correlation of 0.975 and perfect scores of 1.0 in both Spearman and Kendall correlations.
+Additionally, we re-evaluate long-form summarization performance based on EVA-Score.
+The results highlight the impressive capability of GPT-4, especially as the context length increases, where it ranks first on two test datasets and second on the remaining two.
+We also analyze why certain LLMs fail on specific datasets and provide insights into scenarios where EVA-Score aligns with human annotations versus cases where discrepancies occur.
+These analyses shed light on the future development of automatic evaluation metrics for abstractive long-form summarization.
 2 Related Work
-2.1 Long-form Summarization
-Long-form summarization requires language models to understand the purpose of the context and capture core information from a long context which is a struggle for language models.
-During the last few years, much work from different fields regarding long-form summarization has been studied such as news reports (See et al., 2017), government reports (Cohan et al., 2018), academic papers and books (Kryściński et al., 2022), posing great challenges for language models (Lewis et al., 2019; Raffel et al., 2023).
-Since ChatGPT and GPT-4 (OpenAI et al., 2024) come out, LLMs show great potential to deal with long-form summarization (Lu et al., 2023; Zhang et al., 2023a; Wang et al., 2023a) because of their huge amount of pre-training corpus.
-However, when the context length becomes much larger, LLMs still have trouble because of limitations for input sequences and knowledge forgetting when reading a long context.
-More recently, some research aims at exploring LLMs which can handle longer context paving the way for long-form summarization.
-(Su et al., 2023; Li* et al., 2023) Yet, summarizing report-level or book-level texts is still a difficult task.
-(Chang et al., 2024)
+2.1 Abstractive Long-form Summarization
+With the advent of models like ChatGPT and GPT-4 (OpenAI et al., 2024), LLMs have demonstrated considerable potential in traditional summarization tasks (Lu et al., 2023; Zhang et al., 2023; Wang et al., 2023a).
+However, as context length grows, LLMs still face significant issues (Bai et al., 2023).
+Recent research has explored methods to extend the capability of LLMs to handle longer contexts, which meanwhile facilitates long-form summarization (Su et al., 2023; Li* et al., 2023).
+Long-form summarization requires language models to understand the purpose of the context and paraphrases core information from extensive text which can be categorized into extractive and abstractive summarization, with the latter providing a more fluent and flexible summary that better aligns with human preferences compared to extractive methods.
+Therefore, many studies focus on abstractive long-form summarization in different domains, including news reports (See et al., 2017), government reports (Cohan et al., 2018), and books (Kryściński et al., 2022).
+Despite these efforts, abstractive summarization for book-level and report-level texts still remains a challenging task (Chang et al., 2024).
 2.2 Evaluation of Summarization
-Tradition evaluation metrics for summarization include ROUGE variants, like ROUGE-k and ROUGE-L and BERTScore (Zhang et al., 2020).
-However, they only reflect the similarity between candidates and references without the information contained in the prediction which is not enough for us to choose a better summary or quantize the informativeness.
-Currently, several evaluation metrics (Zheng et al., 2024; Ke et al., 2023) based on LLMs have been proposed devoted to aligning model evaluation with human evaluation.
+Tradition evaluation metrics for summarization include ROUGE and its variants, and BERTScore (Zhang et al., 2020).
+However, (Krishna et al., 2021) highlights that ROUGE is not such an informative metric, implying its drawback for abstractive long-form summarization.
+Similar challenges apply to BERTScore.
+Currently, several LLM-based evaluation metrics (Zheng et al., 2024; Ke et al., 2023) have been proposed to aligning model evaluation with human evaluation.
 Some works apply LLMs to the evaluation of long-form summarization, for example, Luo et al.
-(2023) uses ChatGPT to evaluate factual inconsistency for text summarization.
-More Recently, (Jung et al., 2024) calculates the saliency of a summary using MSM(Masked Language Model) to find how a candidate summary can be recovered given a reference summary.
-However, we claim that making use of the inherent ability of LLMs only is not enough to measure the ability of models to summarize for the lack of evaluation of the degree of information contained.
-Moreover, leveraging pre-trained probability distribution is not good enough because of its unexplainability.
-We hope to find a reasonable and explainable method for evaluation.
+(2023) uses ChatGPT to evaluate factual inconsistency for text summarization and Song et al.
+(2024) leverages LLMs to evaluate summaries based on faithfulness, completeness, and conciseness.
+However, we argue that only using the inherent ability of LLMs is not enough to measure the quality of a summary due to the inferior ability of LLMs to differentiate between summaries with nuanced information.
+We are seeking to find a reasonable yet explainable method for the evaluation of long-form summarization.
 3 Preliminaries
-In this section, we provide definitions for elements used in our evaluation pipeline.
-Our work begins with extracting atomic facts from documents.
-Specifically, we denote reference summary as $s$ and candidate summary as $\hat{s}$.
-The extracted atomic fact lists are $S$ and $\hat{S}$ with $s_{i}$ and $\hat{s_{i}}$ being single atomic facts in the list.
-After atomic fact chain generation, we obtain lists containing logic chains $C$ and $\hat{C}$ for reference and candidate, while the elements of them are defined as $C_{i}$ and $\hat{C_{i}}$ and atomic fact in logic chains are defined as $C_{i_{j}}$ and ${\hat{C_{i_{j}}}}$.
-Finally, we conduct Document-level Relation Extraction(DocRE) to complement the original atomic fact list with hierarchical information.
-The final lists are $DocS$ and $Doc\hat{S}$ separately.
+In this section, we provide definitions for terms used in our evaluation pipeline.
+Our work begins with extracting atomic facts from the reference and the candidate which we denote as $Sum$ and as $\hat{Sum}$ respectively.
+The extracted atomic fact lists are $S$ and $\hat{S}$ with $s_{i}$ and $\hat{s_{i}}$ being each atomic fact in the list.
+After Atomic Fact Chain Generation, we obtain fact chain lists $\hat{C}$ from $\hat{S}$, while each element in it is defined $\hat{C_{i}}$ and each atomic fact in logic chains is defined as ${\hat{C_{i_{j}}}}$.
+Finally, we conduct Document-level Relation Extraction to complement the original atomic fact list with hierarchical information.
+The final lists are $DocS$ and $Doc\hat{S}$ separately and the elements in them are denoted as $D_{i}$ and ${\hat{D_{i_{j}}}}$.
 4 Methods
-In this section, we introduce our proposed pipeline to evaluate informativeness in long-form summarization by extraction and validation.
-Given $s$ and $\hat{s}$, our automatic evaluation metric operates by 1) obtaining all atomic facts from $s$ and $\hat{s}$; 2) reforming atomic facts to form logic chains; 3) extracting all document-level relations and filter the duplicated one; 4) using LLM to validate whether it is true or false.
-Figure 3: Examples show the generation of atomic facts. It uses demonstration to help LLMs understand the task.
+In this section, we introduce the detailed implementation of EVA-Score to evaluate informativeness in long-form summarization.
 4.1 Atomic Fact Generation
-Atomic facts are defined as concise sentences that convey a single piece of information.
-In any given sentence within a model prediction, there may be multiple pieces of information that require individual verification.
-Consequently, evaluating the truthfulness of sentences on a one-by-one basis renders the true-or-false outcomes inadequate for assessing informativeness.
-Previous studies find that splitting candidate prediction into atomic facts (Min et al., 2023) improves the evaluation by a large margin and enables a more fine-grained evaluation.
-Following their work, we use ChatGPT with same prompt to generate atomic facts from both $S$ and $\hat{S}$ as $\left[s_{1},s_{2},s_{3}...\right]$ and $\left[\hat{s_{1}},\hat{s_{2}},\hat{s_{3}}...\right]$ where $s_{i}$ is atomic fact from reference summarization and $\hat{s_{i}}$ is atomic fact from candidate summarization.
-4.2 Atomic Fact Chain Generation
-Although claimed atomic, an examination of the “atomic facts” presented reveals that they often encompass more than one atomic fact.
-As illustrated in Figure 3, certain information cannot be independently extracted due to its reliance on contextual semantics.
-Consequently, in cases of cascaded atomic facts, where subsequent facts incorporate or entirely encompass information from preceding facts, if any of the preceding facts are incorrect, all subsequent facts are deemed incorrect, even if the newly added information is accurate.
-To clarify, consider the examples: He played for Berekum Arsenal” and He played for Berekum Arsenal before moving to the United States.” These are two cascaded facts, with the latter not being atomic.
-However, the latter sentence cannot be decomposed into atomic facts because the adverbial clause depends on the entirety of the sentence for its meaning.
-Consequently, if the initial statement is incorrect, any subsequent statements, including those with additional correct information, are also considered incorrect.
-We propose using LLMs, e.g.
-Mixtral-7B-instruct in our work to find logical relationships between atomic facts and treating atomic facts as a list of logic chains.
-The reformed lists are represented as $[C_{1},C_{2},C_{3}...]$ and $[\hat{C_{1}},\hat{C_{2}},\hat{C_{3}}...]$.
+In any given sentence within a model prediction, multiple pieces of information may coexist.
+Consequently, evaluating sentences as a single unit renders the true-or-false outcomes, which are inadequate for assessing the quality.
+Min et al.
+(2023) finds that splitting the prediction into atomic facts, defined as concise sentences consisting of a single piece of information, enables a more fine-grained evaluation and improves the evaluation by a large margin.
+Following their work, we use ChatGPT with the same prompt to generate atomic facts from both $Sum$ and $\hat{Sum}$ and obtain $S$ and $\hat{S}$.
+We share an example in Appendix A.1.
+The prompt used for Atomic Fact Generation is displayed in Appendix C.1.
+The temperature is set to $0.0$ to ensure the effectiveness of demonstrations.
+4.2 Atomic Fact Chain Generation (AFCG)
+Although claimed atomic, certain information, like time and location, cannot be independently extracted due to its reliance on contextual semantics.
+Consequently, parts of $\hat{S}$ may appear as cascaded atomic fact lists, where subsequent facts incorporate or entirely encompass information from preceding facts.
+In this case, if any of the preceding facts are incorrect, all subsequent facts are deemed wrong, even if the newly added information is accurate.
+Since we try to measure the quality of a summary from the information perspective, the former incorrect information should not affect the latter.
+To mitigate the impact, we reform the atomic facts into several logic chains so that in one logic chain, when analyzing a certain atomic fact, we can explicitly focus on newly added information.
+We show an example of AFCG in Appendix A.2.
+In particular, AFCG applies only to $\hat{S}$ since all the information in $Sum$ is considered ground truth.
+Therefore, the cascade effect does not exist.
+We use Mistral-7B-instruct (Jiang et al., 2023a) in our work to find logical relationships between atomic facts, treating it as an NLI task between one fact and the exact next fact, due to the special structure of $\hat{S}$ (Min et al., 2023), and reforming atomic facts as a list of logic chains.
+The prompt used for AFCG is listed in Appendix C.2.
+The temperature is set at $1.0$ to encourage more divergent responses.
 4.3 Document-level Relation Extraction
-Extracting atomic facts from text primarily captures sentence-level relations.
-However, it is crucial to also consider document-level relations, which are challenging to discern solely through atomic fact generation.
-These relations are integral to understanding the contextual logic of a document.
-We complement Document-level relation extractions(DocRE) using the state-of-the-art model GPT-4 with in-context learning for both $s$ and $\hat{s}$ and the results are $[(h_{1},r_{1},t_{1}),(h_{2},r_{2},t_{2})...]$ and $[(\hat{h_{1}},\hat{r_{1}},\hat{t_{1}}),(\hat{h_{2}},\hat{t_{2}},\hat{r_{2}})...]$ where $(h_{i},r_{i},t_{i})$ and $(\hat{h_{i}},\hat{r_{i}},\hat{t_{i}})$ are head entity, relation and tail entity correspondingly.
-We then paraphrase the triples to make them natural language for post-processing.
-To ensure that no sentence-level relations are included and contaminate work later, we use BERTScore along with TF-IDF to filter out the similar ones existing both in the atomic fact list and document-level relation list.
+Extracting atomic facts from text mainly captures sentence-level relations—those between entities within the same sentence.
+However, document-level relations, involving entities across an entire document or beyond sentence boundaries, are crucial for understanding contextual logic.
+To address this, we implement Document-level Relation Extraction (DocRE) using GPT-4, starting with Named Entity Recognition (NER) followed by in-context learning to extract document-level relations from $Sum$ and $\hat{Sum}$.
+The outputs are tuples $[(h_{1},r_{1},t_{1}),(h_{2},r_{2},t_{2}),\dots]$ and $[(\hat{h_{1}},\hat{r_{1}},\hat{t_{1}}),(\hat{h_{2}},\hat{r_{2}},\hat{t_{2}}),\dots]$, representing head entities, relations, and tail entities respectively.
+The temperature for NER and DocRE is set at $1.0$ to ensure diverse outputs.
+Appendix A.3 provides an example illustrating differences between sentence-level and document-level relations, and Appendices 14 and C.4 show the prompts for NER and DocRE, respectively.
+Finally, the triples are paraphrased into natural language for post-processing.
+To prevent the inclusion of sentence-level relations, we use cosine similarity based on BERT (Devlin et al., 2019) to filter out overlapping information between the atomic fact list and the document-level relation list.
+To determine the threshold, we randomly sample a subset from the datasets listed in Section 5.1, test various thresholds, and select the optimal one, as detailed in Appendix B.2.
+We ultimately choose a threshold of $0.65$.
+The filtered relations are then incorporated into $C$ and $\hat{C}$ to form $DocS$ and $\hat{DocS}$ where $DocS$ includes atomic facts from Section 4.1 and document-level relations, whereas $\hat{DocS}$ comprises atomic fact chains from Section 4.2 and document-level relations with each relation as a single-element list.
 4.4 LLM for Validation
-For each atomic fact $s_{i}$ in the candidate atomic fact list, we retrieve $k$ most relevant atomic facts from the reference summarization atomic fact list where k is a hyper-parameter which is a trade-off between accuracy and efficiency.
-For each atomic fact $\hat{C_{i_{j}}}$ in the corresponding logic chain $\hat{C_{i}}$, we check out one by one according to $s_{i}$ from $k$ most relevant atomic facts in $DocS$ represented as $f(\hat{C_{i_{j}}},s_{i})$.
-Specifically, we use Mixtral-7B-Instruct as our validation model.
-By providing the validation model with previous states and prompting it not to pay attention to previous information, we ensure that one piece of information is treated at a time.
+For each atomic fact $\hat{D_{i_{j}}}$ in $\hat{DocS}$, we retrieve $k$ most relevant atomic facts according to BERTScore (Zhang et al., 2020) from $DocS$, using $\hat{D_{i_{j}}}$ as hypothesis and each $D_{i}$ as the reference.
+$k$ is a hyperparameter which is a trade-off between effectiveness and efficiency, and we use $k=3$ in our experiment settings.
+We use Mistral-7B-instruct as our validation model with a temperature equal to $1.0$.
+The matching process between two statements is represented using $f(\cdot)$.
+By providing the correctness of previous statements and prompting it not to pay attention to them but to focus on newly added information, we ensure that one piece of information is checked each time.
+We provide an example in Appendix A.4 and the prompt used for LLMs is listed in Appendix C.5.
 We calculate our Precision and Recall as follows:
-$\text{Precision}=\frac{1}{|Doc\hat{S}|}\sum_{i=0}^{n}\mathds{1}\left(f(\hat{C_{i_{j}}},s_{i})\right)$ (1)
-$\text{Recall}=\frac{1}{|DocS|}\sum_{i=0}^{n}\mathds{1}\left(f^{-1}(s_{i},\hat{C_{i_{j}}})\right)$ (2)
+$\text{Precision}=\frac{1}{|Doc\hat{S}|}\sum_{i=0}^{n}\mathds{1}\left(f(\hat{D_{i_{j}}},D_{i})\right)$ (1)
+$\text{Recall}=\frac{1}{|DocS|}\sum_{i=0}^{n}\mathds{1}\left(f^{-1}(D_{i},\hat{D_{i_{j}}})\right)$ (2)
 where $n$ is the number of atomic facts in $Doc\hat{S}$. We define EVA-Score as the harmonic mean of Precision and Recall.
 5 Experiment
+In this section, we detail our implementations and the results of the experiments.
 5.1 Dataset
-To test the applicability of EVA-Score, we collect long-form summarization with different context lengths.
-Specifically, we adopt CNN Dailymail(See et al., 2017), PubMed, GovReport(Cohan et al., 2018), arXiv, BookSum(Kryściński et al., 2022).
-In particular, we use the chapter-level summarization version of BookSum instead of the book level for less difficulty.
-For all these datasets, we use validation datasets for evaluation.
-Dataset CNN Dailymail Pubmed arXiv Gov Report BookSum
-Average Text Length 781 3111 5894 7797 14337
-Average Summary Length 56 216 172 573 570
-Table 1: The Average Text Length and the Average Summary Length in our experimented dataset.
-5.2 Model
-To perform a fair comparison, we experiment on a wide range of Generative Language Models of different numbers of parameters, and from different architectures.
-For open-source models, we test EVA-Score on Mistral-7B(Jiang et al., 2023) which is a MoE model, variants of Llama-2-chat(Touvron et al., 2023), with parameters 7B and 13B, variants of Vicuna(Zheng et al., 2023) with parameters 7B and 13B and Gemma-7B-it(Team et al., 2024).
-For close-source models, we experiment on ChatGPT and GPT-4(0125-preview version).
+To evaluate the applicability of EVA-Score, we gather long-form summarization across varying context lengths and domains.
+We adopt CNN Dailymail (See et al., 2017), PubMed, GovReport (Cohan et al., 2018), arXiv, and BookSum (Kryściński et al., 2022).
+In particular, we select the chapter-level summarization version of BookSum for less difficulty.
+For all datasets, we use validation datasets for evaluation.
+We list a detailed description of the dataset in Appendix B.1.
 5.2 Baseline
-We compare the results of EVA-Score with traditional evaluation metrics like ROUGE and BERTScore(Zhang et al., 2020) which are based on text overlap and word embeddings.
-Recently, many works have leveraged LLMs to obtain the informativeness of a given text corresponding to a reference.
-We include Vicuna-13B and Mixtral-7B-Instruct as our open-source LLM evaluators and ChatGPT and GPT-4(0125-preview version) as our close-source LLM evaluators.
-For all the LLMs, we instruct them using the prompt from Table 12, imitating the human annotation process.
+We compare the results of EVA-Score with traditional evaluation metrics, i.e., ROUGE-1, ROUGE-2 (Lin, 2004) and BERTScore (Zhang et al., 2020) which are based on text overlap or word embeddings.
+For LLM-based evaluation metrics, we include ChatGPT and GPT-4-0125-preview as our evaluators since they are always considered one of the strongest LLMs, and we denote these methods as prompt-only metrics.
+Moreover, we validate on fine-tuned LLMs including Auto-J-13B (Li et al., 2023) and Critique-LLM-6B (Ke et al., 2024).We denote these two metrics as fine-tuned metrics.
+For ChatGPT and GPT-4, we instruct them using the prompt from Table 17.
+For all the LLMs, we set temperatures at $1.0$.
 5.3 Human Annotation
-We are committed to developing an evaluation metric that resonates with human judgment.
-To achieve this, we randomly select 30 records from the dataset and assess the informativeness of each record by ourselves.
-Initially, we identify sentence-level and document-level atomic facts, denoted as $S$ and $\hat{S}$, from $s$ and $\hat{s}$, respectively.
-They then determine the overlap between $S$ and $\hat{S}$.
+We aim to develop an evaluation metric that resonates with human judgment.
+To achieve this, we randomly select 50 records from the datasets mentioned in Section 5.1 and assess the informativeness of each record.
+We employ five annotators, including some of the authors, all of whom are either M.S.
+or Ph.D.
+students.
+Initially, the annotators identify atomic facts at the sentence and document level, denoted as $DocS$ and $\hat{DocS}$, from $Sum$ and $\hat{Sum}$.
+They then determine the overlap between $DocS$ and $\hat{DocS}$.
 Redundant information, which appears multiple times, is recorded only once.
-Subsequently, we calculate Precision and Recall for these overlaps and derive the F1 Score as their harmonic mean.
-Furthermore, we ensure that each record is evaluated by at least two annotators to maintain assessment fairness.
+Subsequently, they calculate Precision and Recall for these overlaps and derive the F1 Score as their harmonic mean.
+Furthermore, we ensure that each record is evaluated by at least two annotators to maintain the fairness of the assessment.
+Each annotation costs about $15$ minutes.
+The high time cost implies the complexity and high quality of our annotation.
+The detailed annotation recipe is listed in Appendix B.3.
+After the annotation process, we assessed the quality by measuring the agreement between the annotators.
+We achieve an Intraclass Correlation Coefficient (ICC) of 0.999 using a two-way random-effects model, and a Pearson correlation of 0.814.
+These high correlations demonstrate the robustness and reliability of our annotations.
 5.4 Main Results
@@
 BERT-Score 0.534 0.533 0.413 0.623 0.600 0.333
-Model-based Evaluation Metric
-Mixtral-7B-Instruct 0.097 0.052 0.030 -0.653 -0.400 -0.333
-Vicuna-13B 0.065 -0.040 -0.029 0.367 0 0
-ChatGPT 0.138 0.210 0.162 0.430 0.600 0.333
+Trained Evaluation Metric
+Auto-J-13B 0.285 0.233 0.167 0.885 0.947 0.913
+CritiqueLLM-6B 0.510 0.490 0.353 0.865 0.800 0.667
+Prompt-only Evaluation Metric
+ChatGPT 0.135 0.210 0.162 0.430 0.600 0.333
 GPT-4 0.308 0.329 0.229 0.505 0.600 0.333
-EVA-Score (Ours) 0.710 0.673 0.503 0.975 1.000 1.000
-Table 2: Text-level and system-level Pearson($\sigma$), Spearman($\gamma$), and Kendall($\tau$) correlations with different evaluation metrics. The highest correlation among the methods is bold
+EVA-Score 0.710 0.673 0.503 0.975 1.000 1.000
+Table 1: Text-level and system-level Pearson($\sigma$), Spearman($\gamma$), and Kendall($\tau$) correlations between different evaluation metrics and human annotations.
+The highest correlation among the methods is denoted in bold.
+Our EVA-Score achieves the best performance with a perfect correlation at the System-Level.
 In this section, we report the correlation between each evaluation metric and human annotations.
 We adopt text-level and system-level Pearson ($\sigma$), Spearman ($\gamma$), and Kendall ($\tau$) correlation.
-Text-level correlation is calculated by computing the correlation coefficients between human evaluations and automatic evaluation metrics and then averaging these coefficients for all generated texts per instruction.
-System-level correlation, on the other hand, is determined by computing the correlation coefficients between human evaluations and automatic evaluation metrics, but this time for the average score of each dataset.
-From Table 2, we observe that our EVA-Score outperforms all other model-based and traditional evaluation metrics in achieving the highest correlation with human judgments across all indicators at both levels.
+Text-level correlation is calculated by computing the correlation coefficients between human evaluations and automatic evaluation metrics and then averaging these coefficients for all generated texts.
+System-level correlation, on the other hand, is determined by computing the correlation coefficients between human evaluations and automatic evaluation metrics for the average score of each dataset.
+The results are listed in Table 1.
+From the results, we observe that our EVA-Score outperforms all other traditional and LLM-based evaluation metrics in achieving the highest correlation with human judgments across all indicators at both levels.
 At the System Level, EVA-Score achieves perfect Spearman and Kendall correlations of 1.00, demonstrating complete alignment with human assessments.
 In contrast, Traditional Evaluation Methods, while somewhat correlated with human evaluations, generally lack strong correlations, highlighting their limitations.
-Surprisingly, we find that the information scores assigned by Vicuna-13B show little correlation with humans, and At the System Level, the scores from Mistral-7B show even negative correlation with humans.
-We attribute this to the situation that the informativeness is obtained by quantitative calculation using EVA-Score and Human Annotation.
-The instruction may be too complex for them to deal with within a single interaction which requires them first to extract atomic facts and document-level relations and verify all of them one by one.
-However, for models like ChatGPT and GPT-4, while their correlation with human judgments at the Text Level remains minimal, they exhibit promising outcomes at the System Level using only In-Context Learning.
-We also report our results on several long-form summarization datasets, including CNN Dailymail, PubMed, GovReport, arXiv, and BookSum, using three different evaluation metrics, i.e.
-EVA-Score, ROUGE-1, and BERTScore.
-Since this work focuses on the evaluation of long-form summarization, we don’t pursue feeding long context into LLMs without information loss.
-Instead, we believe that useful information mostly appears at the beginning or the end of a text, we just simply truncate the context from both the beginning and end to meet the models’ max input sequence length.
+Models such as ChatGPT and GPT-4 exhibit a relatively low correlation with human annotations, which we attribute to the complexity involved in assessing the informativeness of a given summary against a reference, requiring them to decompose the task and finish the task step by step.
+Interestingly, evaluation models fine-tuned specifically for NLG achieve a high correlation with human annotations at the System Level.
+This improved performance is likely due to their fine-tuning process, which explicitly includes informativeness as an evaluation criterion, suggesting a promising direction for the future development of larger evaluation models.
+5.5 Error Analysis
 Dataset CNN Dailymail Pubmed arXiv Gov Report BookSum
@@
 Vicuna-13B 0.462 0.432 0.872 0.431 0.388 0.837 0.346 0.348 0.832 0.154 0.298 0.849 0.144 0.188 0.827
-Mixtral-7B-Instruct 0.499 0.422 0.872 0.447 0.447 0.847 0.444 0.448 0.845 0.313 0.399 0.854 0.259 0.377 0.839
+Mistral-7B-Instruct 0.499 0.422 0.872 0.447 0.447 0.847 0.444 0.448 0.845 0.313 0.399 0.854 0.259 0.377 0.839
 Llama-2-7B-chat 0.399 0.386 0.868 0.358 0.421 0.845 0.411 0.435 0.851 0.193 0.198 0.798 0.186 0.268 0.826
@@
 Table 2: The results of several LLMs on different long-form summarization datasets.
-We denote EVA-Score as EVA, R-1 as ROUGE-1, BS as BERTScore.
+We denote EVA-Score as EVA, R-1 as ROUGE-1, and BS as BERTScore.
 All the values are normalized between 0 and 1.
 We indicate the highest score of EVA-Score using bold.
-Referring to Table 3, we observe that as the context lengthens, all tested models experience varying degrees of performance decline.
-Particularly notable is the performance at the part-of-book level, i.e.
-chapter summarization in BookSum, where the results are markedly poor compared to the ground truth.
-This issue, however, can be mitigated by employing methods that feed the entire context into the model, rather than truncating it as we have done in our initial experiments.
-Additionally, compared to other evaluation metrics, the EVA-Score demonstrates a more pronounced and consistent decline in performance as context length increases, with fewer fluctuations.
-This pattern aligns well with human cognitive processes, indicating that as information load increases, the challenge of maintaining high summarization quality also grows.
-This consistent pattern of decline not only reflects realistic expectations of model performance but also underscores the effectiveness of the EVA-Score in capturing these dynamics, further validating its utility from another perspective.
-Moreover, models using ROPE(Su et al., 2023) or other methods, enabling longer input sequences show greater performance in long-form summarization when the context length becomes larger, mainly due to their ability to comprehend a long context or search in a long context.
-However, when the context is not so long, they show comparable or inferior performance with assistant models with normal sequence length, e.g.
-4096 or 8192.
-Under the evaluation framework of the EVA-Score, GPT-4 demonstrates robust capabilities, particularly with longer input sequences, compared to other models.
-In the CNN Dailymail dataset, it trails behind models such as Vicuna-7B and ChatGPT.
-This is primarily due to GPT-4’s tendency to generate more extensive summaries than its counterparts in settings with relatively short contexts, thereby increasing its count of atomic facts.
-While this might allow it to cover more points than other models, it often results in lower precision due to the inclusion of information that extends beyond the reference summary.
-Nevertheless, in other datasets such as arXiv and Gov Report, GPT-4 outperforms all competitors.
-It also secures a second-place ranking in Pubmed and BookSum, affirming its effectiveness across diverse text summarization tasks.
-To be mentioned, Mixtral-7B-Instruct, compared with other models with the same parameter or even larger, shows great potential to fulfill long-form summarization and even beats GPT-4 by a small margin in chapter-level BookSum, which meets the findings of Chang et al.
+To have a thorough understanding of EVA-Score, we provide error analysis from the pipeline of EVA-Score in Appendix B.5 showing several situations where EVA-Score fails.
+To step further into why ROUGE and BERTScore are more suitable in relatively shorter settings than in longer settings, we provide our explanation in Appendix B.6.
+5.6 Re-evaluation on Abstractive Long-form Summarization
+To have a thorough understanding of the performance of LLMs on abstractive long-form summarization, we test on a wide range of models.
+For open source models, we evaluate on Mistral-7B-instruct (Jiang et al., 2023b), Llama-2-chat-7B, Llama-2-chat-13B (Touvron et al., 2023), Vicuna-7B, Vicuna-13B (Zheng et al., 2023), and Gemma-7B-it (Team et al., 2024).
+For close-source models, we experiment with ChatGPT and GPT-4-0125-preview.
+For all the LLMs, we set the temperatures as $1.0$.
+We use three different evaluation metrics to indicate their performance, i.e.
+EVA-Score, ROUGE-1, and BERTScore.
+Since this work focuses on the evaluation of long-form summarization, we do not pursue feeding long context into LLMs without information loss.
+Instead, we believe that useful information usually appears at the beginning or end of a text as Bai et al.
+(2023), and we simply truncate the context from the beginning and end to meet the maximum input sequence length of the models.
+The length of the context window can be viewed as part of the models’ inherent abilities, so it will not introduce unfairness in our evaluation.
+For all GLMs, our prompt is “Summary the following text.
+You can use your own words instead of extracting from the original document.”, ensuring an abstractive setting.
+Referring to Table 2, as context length increases, all models experience varying degrees of performance decline, particularly in the BookSum dataset, where results are notably poorer.
+However, models capable of handling longer input sequences (Su et al., 2023) demonstrate superior performance in abstractive long-form summarization compared to those with limited context windows when dealing with larger contexts, largely due to their ability to comprehend extensive content.
+Conversely, when the context length is shorter, these models perform comparably or even worse than models with normal sequence lengths, such as Vicuna-7B.
+GPT-4 demonstrates robust capabilities compared to others, particularly in situations of longer input sequences under EVA-Score.
+However, it fails in relatively shorter settings, probably resulting from the tendency to respond in more words than others.
+Surprisingly, Mistral-7B-instruct, compared with other models with the same amount of parameters or even more parameters, shows great potential to fulfill abstractive long-form summarization and even beats GPT-4 by a small margin in BookSum, which meets the findings of Chang et al.
 (2024).
+Moreover, compared to other evaluation metrics, EVA-Score shows a more pronounced and consistent decrease in performance as the context length grows, with fewer fluctuations.
+This pattern mirrors human cognitive processes, where increasing information load makes maintaining summarization quality more challenging.
+The consistent decline highlights the effectiveness of EVA-Score from another perspective.
 5.7 Analysis of Summarization Length
 Dataset CNN Dailymail Pubmed arXiv Gov Report Booksum
-Indicator Avg Length Avg Length Avg Length Avg Length Avg Length
+Metric Avg Length Avg Length Avg Length Avg Length Avg Length
 Vicuna-7B 108 214 133 258 124
@@
 Mistral-7B-Instruct 109 207 212 353 251
-Llama-2-7B-chat-hf 118 143 108 140 185
-Llama-2-13B-chat-hf 124 184 221 236 179
+Llama-2-7B-chat 118 143 108 140 185
+Llama-2-13B-chat 124 184 221 236 179
 Gemma-7B-it 98 139 140 177 151
@@
 GPT-4 153 227 272 242 212
-Table 4: The average summarization length(denoted by Avg Length by different model. The longest is marked underline
-To elucidate why GPT-4 underperforms on the EVA-Score with a shorter context, we conduct an analysis focused on the summarization lengths produced by different models, as detailed in Table 4.
-Models pre-trained with longer input sequences tend to generate more extensive responses.
-Specifically, GPT-4 produces summaries of 153 tokens in CNN Dailymail, 227 tokens in Pubmed, and 272 tokens in arXiv and Mixtral-7B-Instruct outputs 353 tokens in Gov Report and 251 tokens in BookSum whose maximum input sequences length are 128000 and 32768 respectively.
-Upon manual examination, we observe that shorter contexts typically result in shorter ground truth.
-Consequently, if a model generates a lengthy prediction, this verbosity is penalized with a reduced EVA-Score due to diminished precision.
-As the context lengthens, references become richer in information.
-However, these models tend to maintain a similar response length, lacking significant information addition, thereby failing to capture enough relevant points, especially in longer predictions.
-This discrepancy explains GPT-4’s shortfall in CNN Dailymail and the superior performance of Mistral-7B-Instruct, despite having a comparable number of parameters.
-Furthermore, GPT-4, which is refined through Reinforcement Learning from Human Feedback (RLHF), typically caps its responses to summarizations under 300 tokens, even in contexts of substantial length.
-This behavior is a contributing factor to GPT-4’s underachievement in BookSum, where more extensive and detailed responses are crucial for capturing the full scope of the content.
+Table 3: The average length of summary counted in words generated by different models denoted by Avg Length.
+The largest count is marked using bold.
+It can be seen that GPT-4 prefers to summarize with the length of responses in a narrow range.
+To elucidate why GPT-4 underperforms under EVA-Score in the situation of a shorter context, we analyze the summary lengths produced by different models, as detailed in Table 3.
+We observe that shorter contexts typically result in shorter ground truth.
+Consequently, if a model generates a lengthy summary, this verbosity is penalized with a reduced EVA-Score due to decreased precision.
+As the context lengthens, the references become richer in information.
+If these models tend to maintain a similar response length, lacking significant information addition, they fail to capture enough relevant points.
+GPT-4, which is refined through RLHF (Ouyang et al., 2022), typically caps its responses to summarizations to a narrow range.
+This behavior contributes to GPT-4’s underachievement in Booksum, where more extensive and detailed responses are crucial to capture the full scope of the content.
+Additionally, we provide a meta-evaluation of human annotations in Appendix B.4, highlighting the underlying approach adopted by humans and models for summarization.
 5.8 Ablation Study
-We report the effect of Atomic Fact Chain Generation and Document-level Relation Extraction on the number of atomic facts to be verified.
+We report the effect of AFCG and DocRE on the number of atomic facts to be verified.
 Settings AFG only AFCG AFCG + DocRE
@@
 BookSum 36 27 31
-Table 5: Number of facts(or fact chains) to be evaluated under three settings, averaging 9 models we used for experiments, i.e.
-Atomic Fact Generation(AFG), Atomic Fact Chain Generation(AFCG) and Atomic Fact Chain Generation Complemented with DocRE(AFCG + DocRE).
-From Table 5, datasets with longer context and richer information tend to result in more facts, and by using Atomic Fact Chain Generation, we reduce the number of cascaded facts a lot.
-By doing so, minor mistakes in one piece of information will not affect others.
+Table 4: Number of facts (fact chains) to be evaluated under three settings, averaging 9 models we used for experiments, i.e., Atomic Fact Generation (AFG), Atomic Fact Chain Generation (AFCG) and Atomic Fact Chain Generation Complemented with Document-level Relation Extraction (AFCG + DocRE).
+From Table 4, datasets with longer context and richer information tend to result in more facts, and by AFCG, we largely reduce the number of cascaded facts.
 Also, DocRE complements AFG with document-level relations, which is helpful for the evaluation of the hierarchical information in the summary.
-5.8 Meta Evaluation of Human Annotation
-By observing the results of human annotations, we find that ground truth prefers to summarize more elegantly, using core arguments from the original context in addition to examples supported with other a little bit trivial arguments.
+The two processes work together to form a concise yet comprehensive list of facts to be verified.
+5.9 Analysis of Different LLM Verifier
+Since our LLM Verifier should consider previous states in a logic chain and verify the accuracy of current atomic fact, it is a rather hard problem compared to simple verification of an NLI problem.
+Intuitively, with stronger models, we will get a more human-aligned evaluation metric.
+Therefore, we report the results using ChatGPT for a simple proof which are listed in Table 5.
+Leveraging ChatGPT results in a higher human correlation in Pearson and Kendall correlation and comparable Spearman correlation, which is consistent with our recognition.
+However, using ChatGPT results in much more costs than open-source LLMs because of the great number of validations.
+Metric $\sigma$ $\gamma$ $\tau$
+EVA-Score-Mixtral-7B-Instruct 0.710 0.673 0.503
+EVA-Score-ChatGPT 0.737 0.672 0.543
+Table 5: Human correlation using ChatGPT as LLM verifier of EVA-Score compared with EVA-Score with Mistral-7B-Instruct.
+$\sigma$, $\gamma$, $\tau$ are Pearson, Spearman, and Kendall correlation respectively.
+The larger one is denoted in bold.
+5.10 Case Study
+To provide a more intuitive and straightforward insight into the pipeline of EVA-Score, we provide a detailed example in Appendix A.5.
+From the case, we can see that EVA-Score provides not only a quantitative and objective information score but also provides a human-like explanation.
+6 Conclusion and Future Work
+In this work, we introduce a novel evaluation metric, termed EVA-Score, to evaluate the informativeness in abstract long-form summarization.
+EVA-Score begins by extracting atomic facts from given summaries, which are then restructured into logical chains.
+We also extract document-level relations for the completion of information.
+Subsequently, LLMs are employed for further validation of the information.
+Experimental results demonstrate that, compared to existing automatic evaluation metrics, EVA-Score achieves the highest correlation with human judgments, providing an objective and explainable assessment.
+Furthermore, we use EVA-Score to re-evaluate abstractive long-form summarization models, highlighting the challenges LLMs face in handling extended contexts.
+We also conduct a detailed analysis of cases where EVA-Score and other evaluation metrics fall short in evaluating abstractive long-form summaries.
+We believe that EVA-Score, designed to quantify information overlap between reference and candidate, has potential applications beyond long-form summarization and can be extended to broader evaluation scenarios.
+Limitations
+EVA-Score is time-consuming to run
+EVA-Score includes several pre-processing methods.
+We should first extract atomic facts and document-level relations from two summaries.
+We also use LLMs for paraphrasing to make the relations more natural.
+The preprocessing takes around two minutes per record.
+Then for each atomic fact in the candidate, we select $k$ most relevant from the reference and use LLMs for validation.
+Since the number of atomic facts is very large, it takes quite a lot of time.
+We report the time consumed correlated to the sum of the length of the candidate and reference counted in tokens.
+Length Interval 0-300 300-600 600-900 900-1200
+Time Consuming/s 195 402 578 629
+Table 6: Time consumed measured in seconds for different summary lengths measured in tokens. The time consumption shows an approximately linear correlation with text length.
+From Table 6, for a summary of relatively short length, it takes about three minutes to finish the evaluation, while for a longer summary, it may take about ten minutes which is a long time compared to traditional methods.
+EVA-Score only considers Informativeness
+EVA-Score is designed to evaluate informativeness.
+Some existing evaluation metrics consider many other factors, like faithfulness, and consistency.
+EVA-Score may be combined with other evaluation metrics to provide a more comprehensive and fair evaluation of abstractive long-form summarization.
+EVA-Score Leverages GPT-4 for Doc-RE
+EVA-Score incorporates DocRE to address the lack of hierarchical information.
+However, DocRE by GPT-4 may be incomplete, as GPT-4 may struggle to fully capture the core meaning at the document level due to the inherent complexity of the task.
+Leveraging more advanced models specifically fine-tuned for DocRE could potentially improve results.
+Furthermore, the filter used in the current approach is simplistic, and experimental findings suggest that a more effective filter is needed to better distinguish document-level relations from sentence-level ones.
+Acknowledgments
+This work is supported by the National Science and Technology Major Project(2023ZD0121403).
+We express our gratitude to the anonymous reviewers for their insightful feedback.
+We thank Chen Ling for the discussion with us.
+Appendix A Case Study
+In this section, we list all all the cases for display and comprehension.
+A.1 Atomic Fact Generation (AFG)
+We display an example of our AFG process in Figure 3. It decomposes a text into a combination of multiple pieces of information.
+Figure 3: Examples show the generation of atomic facts. We implement it using In-Context-Learning for better extraction.
+A.2 Atomic Fact Chain Generation (AFCG)
+We display an example of our AFCG process in Figure 4.
+We convert the original 8 atomic facts into 2 fact chains to ease validation because each fact in a certain fact chain shares some information overlap.
+Also, by AFCG, we ensure checking pieces of information one by one is feasible.
+Figure 4: An example displaying the process of AFCG.
+We first extract atomic facts from a given summary using AFG.
+Then based on the logic overlap, we reformulate them into several logic chains.
+The texts denoted in red and blue are information cascaded which need to be verified later.
+A.3 Document-level Relation Extraction (DocRE)
+We display an example illustrating the difference between document-level relations and sentence-level relations in Table 7.
+To make it clear, sentence-level relations are relations that exist within a sentence while document-level relations are relations beyond a single sentence that can be found via implicit or explicit reasoning between sentence-level relations.
+Text:
+TechCorp, headquartered in Silicon Valley, is known for its innovative products in the tech industry.
+John recently joined TechCorp as a software engineer.
+He is excited to work on cutting-edge AI technologies.
+Sentence-level Relation:
+(John, employed by, TechCorp)
+Document-level Relations:
+(John, lived in, Silicon Valley)
+Table 7: An example indicating the difference between sentence-level relations and document-level relations.
+A.4 LLM Validation
+We provide an example revealing the process of using LLMs for final validation in Figure 5.
+This validation process verifies newly added information only rather than treating the whole sentence as a basic unit.
+Therefore, we can measure the informativeness from a fine-grained dimension.
+Dataset CNN Dailymail Pubmed arXiv Gov Report BookSum
+Average Text Length 781 3111 5894 7797 14337
+Average Summary Length 56 216 172 573 570
+Table 8: The Average Text Length and the Average Summary Length counted in words in our experimented dataset. The largest values are denoted using bold.
+Figure 5: An example displaying the process of LLM Validation.
+For each fact chain illustrated in Figure 4, we verify the atomic fact within it one by one by explicitly prompting LLMs to focus on newly added information rather than redundant information.
+A.5 Overall Process
+To provide a more intuitive and straightforward insight into the pipeline of EVA-Score, we provide a detailed case in Table 9.
+From the case, we can see that EVA-Score provides not only a quantitative and objective information score but also provides a human-like explanation.
+Appendix B Detailed Experiments
+In this section, we provide details that are not listed in the main papers due to the page limitation.
+B.1 Dataset Details
+We provide the average text length and the average gold summary length, both measured in words, in Table 8.
+Among the datasets, BookSum contains the longest average text length, while Gov Report has the longest average summary length.
+B.2 Threshold for Filtering
+In this section, we explain our approach for selecting the threshold used to filter document-level relations from sentence-level relations.
+Our principle is to filter out as many similar statements as possible while retaining those that are distinct.
+We utilize cosine similarity based on BERT as the filtering criterion.
+Specifically, for each element in the set of extracted document-level relations, we calculate the cosine similarity between it and each fact in the list of extracted atomic facts and then determine if the highest cosine similarity exceeds the threshold.
+To guide our threshold selection, we consider three indicators: (1) Accuracy Rate, which measures the proportion of filtered document-level relations that genuinely required filtering; (2) Remaining Rate, which represents the ratio of document-level relations that are retained after filtering; and (3) Remaining Accuracy Rate, which indicates the proportion of retained document-level relations that genuinely should not have been filtered.
+The results are listed in Table 10.
+From the experimental results, a larger threshold results in fewer statements being filtered out, as only those highly similar to one of the extracted atomic facts are removed.
+This leads to a higher accuracy rate and remaining rate but a lower remaining accuracy rate.
+Based on this trade-off, we set the threshold at 0.65 to minimize redundancy while ensuring sufficient remaining accuracy.
+The Remaining Accuracy Rate is not higher primarily due to the straightforward use of GPT-4 for DocRE.
+However, since the Remaining Rate is lower and there is a general lack of document-level relations within the task, the impact of this limitation is not particularly pronounced due to their little contribution to the final results.
+Reference Summary:
+The mpp decreased the number of dopaminergic neurons by 40% , and increased the release of lactate dehydrogenase ( ldh ) into the culture medium .
+The tq significantly rescued dopaminergic neurons and decreased the release of ldh at the concentrations of 0.1 and 1 m
+Atomic Fact Chains
+[- The mpp decreased dopaminergic neurons, - The mpp decreased the number of dopaminergic neurons, - The mpp decreased the number of dopaminergic neurons by 40%]
+[- The mpp increased the release of lactate dehydrogenase ( ldh ), - The mpp increased the release of lactate dehydrogenase ( ldh ) into the culture medium]
+[- The tq significantly rescued dopaminergic neurons]
+[- The tq decreased the release of ldh, - The tq decreased the release of ldh at the concentrations, - The tq decreased the release of ldh at the concentrations of 0.1 and 1 m]
+Document-level Relations:
+tq causes Thymoquinone
+mpp leads to 1-methyl-4-phenylpyridinium
+Candidate Summary:
+The MPP significantly decreased the survival of dopaminergic neurons and increased the release of lysosomal hydrolase (LDH) into the culture medium.
+The neurotoxicity of MPP was attributed to its selective uptake by dopaminergic neurons through the dopamine transporter.
+Atomic Fact Chains
+[- The MPP significantly decreased dopaminergic neurons, -The MPP significantly decreased the survival of dopaminergic neurons ]
+[- The MPP increased the release of lysosomal hydrolase (LDH), - The MPP increased the release of lysosomal hydrolase (LDH) into the culture medium]
+[- The neurotoxicity was attributed to dopaminergic neurons, - The neurotoxicity of MPP was attributed to dopaminergic neurons, - The neurotoxicity of MPP was attributed to its selective uptake by dopaminergic neurons, - The neurotoxicity of MPP was attributed to dopaminergic neurons through the dopamine transporter, - The neurotoxicity of MPP was attributed to its selective uptake by dopaminergic neurons through the dopamine transporter.]
+Document-level Relations:
+mpp causes 1-methyl-4-phenylpyridinium
+neurotoxicity infects neuroprotective effect
+Matching Process
+- “The mpp decreased dopaminergic neurons” match “The MPP significantly decreased dopaminergic neurons”
+- “The mpp increased the release of lactate dehydrogenase ( ldh )” match “The mpp increased the release of lactate dehydrogenase ( ldh )”
+- “The mpp increased the release of lactate dehydrogenase ( ldh ) into the culture medium” match “The MPP increased the release of lysosomal hydrolase (LDH) into the culture medium”
+- “The mpp decreased dopaminergic neurons and increased the release of lactate dehydrogenase ( ldh )” match ‘The MPP significantly decreased dopaminergic neurons and increased the release of lysosomal hydrolase (LDH)”
+- “The mpp decreased dopaminergic neurons and increased the release of lactate dehydrogenase ( ldh ) into the culture medium” match “The MPP significantly decreased dopaminergic neurons and increased the release of lysosomal hydrolase (LDH) into the culture medium”
+- “mpp causes 1-methyl-4-phenylpyridinium” match “mpp leads to 1-methyl-4-phenylpyridinium”
+Table 9: Case Study For EVA-Score from Pubmed Dataset.
+B.3 Annotation Recipe
+Identify Atomic Information
+We identify information contained in the candidate summary and the reference summary.
+Specifically, for a sentence, there may be several pieces of information so that we can treat it as a combination of information.
+For duplicated information, we only record once to ensure fairness and accuracy.
+Compare Candidate Based on Reference
+To check whether the information contained in the reference also exists in the candidate, we verify information piece by piece.
+A piece of information will only be recorded as overlapping if and only if it exists both in the reference and the candidate.
+After that, we can get Precision and Recall for each record and we use the F1 Score as our final results.
+After Annotation Validation
+To ensure reliability, we ask at least two annotators for annotation for each record.
+After annotation, we calculate the correlation between the annotators.
+If there is an apparent low correlation between annotators, we ask a third annotator for re-annotation.
+B.4 Meta-evaluation of Human Annotation
+By observing the results of human annotations, we find that the ground truth prefers to summarize more elegantly, using core arguments from the original context in addition to examples supported with other a little bit trivial arguments.
 However, LLMs try to list all of the arguments, ignoring the examples or data in the context, leading to differences with ground truth.
 However, there does exist some verbosity in the ground truth, indicating the natural drawback of the summarization dataset.
-5.9 Analysis of Different LLM Verifier
-Since our LLM Verifier should consider previous states in a logic chain and verify whether the current atomic fact is true or not, it is a rather hard problem compared with simply verifying.
-Intuitively, with more capable models, we will get a more human-aligned evaluation metric.
-We report the results using ChatGPT for a simple proof.
-Metric $\sigma$ $\gamma$ $\tau$
-EVA-Score-Mistral-7B-Instruct 0.710 0.673 0.503
-EVA-Score-ChatGPT 0.737 0.672 0.543
-Table 6: Human correlation using ChatGPT as EVA-Score LLM verifier compared with EVA-Score with Mistral-7B-Instruct. The larger one is denoted is bold.
-Leveraging ChatGPT results in more human correlation in Pearson and Kendall correlation and comparable Spearman correlation, which is consistent with our recognition.
-However, using ChatGPT results in much more costs than open-source LLMs because of the great number of validations.
-5.10 Case Study
-To provide a more intuitive and straightforward insight into the pipeline of EVA-Score, we provide a detailed case in Appendix B.
-From the case, we can see that EVA-Score provides not only a quantitative and objective information score but also provides a human-like explanation.
-6 Conclusion and Future Work
-In this work, we propose a new evaluation metric called EVA-Score to evaluate informativeness in long-form summarization.
-It extracts atomic facts which are then reformed as logic chains and document-level relations from reference and candidate and uses LLMs for validation one by one.
-Experiment results show that compared with other traditional metrics or LLM-based metrics, our work shows the best human correlation as well as provides an objective explanation.
-Moreover, we re-evaluate long-form summarization based on EVA-Score, indicating a great challenge in book-level summarization.
-We believe that EVA-Score is meant to calculate information overlap between reference and candidate, thus it can be applied to broader scenarios rather than long-form summarization only.
-In future work, we suggest: 1) Leveraging EVA-Score to evaluate more reference-based and information-rich tasks and 2) Utilizing EVA-Score for the generation of more informative texts.
-Limitations
-EVA-Score can be time-consuming to run
-EVA-Score includes several pre-processing methods.
-We should first extract atomic facts and document-level relations two summaries and use LLMs for paraphrasing to make the relations more natural.
-The pre-processing takes around two minutes per record.
-Then for each atomic fact in the candidate, we select $k$ most relevant from the reference and use LLMs for validation.
-Since the number of atomic facts takes quite a lot of time.
-We report the time consumed correlated to the sum of the length of the candidate and reference.
-Length Interval/token 0-300 300-600 600-900 900-1200
-Time Consuming/s 195 402 578 629
-Table 7: Time consumed regarding to different summary length
-From Table 7, for a summary of relatively short length, it takes about three minutes to finish the evaluation, while for a longer summary, it may take about ten minutes which is a long time compared to traditional methods.
-EVA-Score only considers Informativeness
-Some existing evaluation metrics consider many other factors, like faithfulness, and consistency.
-EVA-Score may be combined with other evaluation metrics to provide a more comprehensive evaluation of long-form summarization.
-EVA-Score Leverages GPT-4 for Doc-RE
-EVA-Score implements Document-level Relation Extraction for the lack of hierarchical information.
-However, Doc-RE by GPT-4 maybe incomplete because GPT-4 may fail to capture the core meaning of the document level due to the difficulty of such a task.
-Leveraging more capable models that are fine-tuned on Doc-RE may be more helpful.
-Acknowledgments
-This work is supported by the National Science and Technology Major Project(2023ZD0121403). We extend our gratitude to the anonymous reviewers for their insightful feedback.
-Appendix A Instructions for EVA-Score
-A.1 Instruction for Atomic Fact Generation
-We use the following prompt to generate atomic facts.
+B.5 Error Analysis
+We provide an error analysis of why EVA-Score underperforms compared to human annotations.
+Since EVA-Score is built on Atomic Fact Generation, Atomic Fact Chain Generation, Document-level Relation Extraction (DocRE), and LLM Validation, we randomly select 30 records to identify the sources of errors.
+Upon manual inspection, we find that most errors occurred during DocRE and LLM Validation.
+This is because AFG is fundamental for LLMs, and AFCG, which involves NLI, is generally handled well by LLMs.
+However, for DocRE, GPT-4 struggles to differentiate between sentence-level and document-level relations, leading to inaccuracies.
+Additionally, the static threshold used for filtering document-level relations may be either too strict or too lenient, further contributing to errors.
+In LLM Validation, we prompt the model to focus on newly added information, which requires a high level of reasoning capability of LLMs.
+We also display examples in Table 11.
+To make EVA-Score more aligned with human judgments, we should either use more powerful LLMs for DocRE or develop novel methods for relation extraction.
+Furthermore, the LLM validation process can be improved through techniques like majority voting or self-consistency, which could enhance the reliability of the evaluations.
+Threshold 0.60 0.65 0.70 0.75
+Accuracy Rate 0.823 0.861 0.886 0.920
+Remaining Rate 0.075 0.142 0.258 0.368
+Remaining Accuracy Rate 0.512 0.476 0.437 0.413
+Table 10: Results of the performance of the filter under different thresholds.
+We use three metrics, Accuracy Rate, Remaining Rate, and Remaining Accuracy Rate, for help.
+The largest value is denoted using bold.
+B.6 Why Traditional Metrics Fail
+The challenge of using traditional metrics in long contexts stems from the fact that surface-level similarity does not always indicate informativeness.
+Metrics like BERTScore and ROUGE measure the similarity between generated text and a reference in terms of content and structure, but they often fall short in assessing the quality and relevance of the conveyed information.
+In long contexts, a text may be highly informative and similar in meaning to the reference, yet exhibit low similarity in terms of word choice or embeddings.
+BERTScore, for instance, faces limitations in longer texts due to semantic dilution, where embeddings capture a wide range of content, including irrelevant details, diluting the importance of key informative elements.
+Its reliance on contextual embeddings also means that the meaning of words can shift with varying contexts.
+Furthermore, as the number of tokens increases, the complexity of token matching may not align with the most informative parts of the text, leading to lower similarity scores despite the presence of valuable information.
+ROUGE, on the other hand, struggles with lexical overlap insufficiency, as it evaluates texts based on overlapping n-grams, failing to account for paraphrases or semantic equivalents.
+This can penalize informative texts that use different wording to convey the same ideas.
+In longer context, the number of possible n-gram overlaps grows, but the proportion of meaningful overlaps decreases, causing the metric to become less sensitive to informativeness.
+Type: Mixture of Doument-level Relations with Sentence-level Relations
+Text:
+The paper presents a theoretical study on the precession of vortex lines in a superfluid in the presence of an external potential.
+The author uses the Gross-Pitaevskii equation to study the stationary states of the system and finds that there are two types of vortex lines that can be described by the lyapunov functional.
+The author also finds that the angular momentum of the vortex line as a function of the precession frequency.
+The paper also compares the results with previous works and finds that the difference in the density of the condensates affects the precession frequency.
+The author concludes that the present results and the density of the condensates.
+Document-level Relation:
+"The theoretical study is on the precession of vortex lines."
+"The Gross-Pitaevskii equation is used to study the stationary states of the system."
+"The author finds that there are two types of vortex lines." $\rightarrow$ It is a sentence-level relation
+"The author finds the angular momentum of the vortex line." $\rightarrow$ It is a sentence-level relation
+"The angular momentum of the vortex line is a function of the precession frequency."
+"Stationary states are a feature of Vortex lines"
+"Angular momentum is a property of vortex lines"
+Type: Miss of Document-level Relations
+Text:
+in that light, it seems surprising that they can say much about inflation at all , but it turns out that the flow equations can be viewed as a ( rather complicated ) algorithm for generating functions which have a suitable form to be interpreted as inflationary model implied that the flow equation predictions ought to be little changed by moving to the braneworld , although this modifies the friedmann equation .
+Doument-level Relations Caught
+"The Friedmann equation was modified in the braneworld inflation scenario"
+Document-level Relations Missed
+"The flow equation predictions little changed in the braneworld inflation scenario"
+Type: LLM Validation Error
+Reference:
+"The scheme uses the two-channel Raman interaction."
+"The interaction is between two atoms."
+"The scheme involves two atoms."
+Hypothesis:
+"The scheme is for implementing an unconventional geometric two-qubit phase gate."
+"The scheme involves using the two-channel Raman interaction."
+"The two atoms are in a cavity."
+Matching Pool:
+"The scheme uses the two-channel Raman interaction." $\rightarrow$ "The scheme involves using the two-channel Raman interaction."
+"The scheme involves two atoms." $\rightarrow$ "The scheme uses the two-channel Raman interaction."
+Table 11: The error analysis of why EVA-Score fails.
+We display the example of Mixture of Document-level Relations with Sentence-level Relations, Miss of Document-level Relations, and LLM Validation Error.
 Instruction: Please breakdown the following sentence into independent facts:
@@
 - He served as the Command Module Pilot for the Gemini 7 mission.
-Table 8: The instruction for atomic fact generation
-A.2 Instructon for Atomic Fact Chain Generation
-As discussed before, we use the following prompt to decide whether one fact is entailed by others.
+Table 12: The instruction for AFG.
 Instruction: Given a hypothesis and a reference, determine if the hypothesis is entailed by the reference.
@@
 Answer: False
-Table 9: The instruction for atomic fact chain generation
-A.3 Instruction for Document-level Relation Extraction
-We use the following prompt with GPT-4 to conduct Document-level Relation Extraction.
+Table 13: The instruction for AFCG.
+Appendix C Instructions for EVA-Score
+C.1 Instruction for Atomic Fact Generation
+We use the prompt in Table 12 to generate atomic facts.
+C.2 Instruction for Atomic Fact Chain Generation
+As discussed before, we use the prompt in Table 13 to decide whether one fact is entailed by others.
+C.3 Instruction for Named Entity Recognition
+We use the prompt in Table 14 to get named entities from the given text.
+Instruction: Given a text, your task is to extract all the entities as comprehensive as possible from the text.
+Below is the list of entities that you need to extract from the text.
+| Types | Content |
+| PER | People, including fictional |
+| ORG | Companies, universities, institutions, political or religious groups, etc. |
+| LOC | Geographically defined locations, including mountains, waters, etc.
+Politically defined locations, including countries, cities, states, streets, etc.
+Facilities, including buildings, museums, stadiums, hospitals, factories, airports, etc.
+|
+| TIME | Absolute or relative dates or periods. |
+| NUM | Percents, money, quantities Products, including vehicles, weapons, etc. |
+| MISC | Events, including elections, battles, sporting events, etc. Laws, cases, languages, etc |
+Respond with a list of dictionaries, where each dictionary contains the entity type and the entity text.
+The list should be sorted by the entity type, and the entities should be sorted by their start position.
+Note: You should respond in a list of dictionaries, where each dictionary contains the entity type and the entity text.
+To make it specific, the format is like: [
+"type": "", "text": "",
+"type": "", "text": "",
+"type": "", "text": ""
+I will provide you with some examples to help you understand the task.
+Domonstrations
+Example 1 "Apple is looking at buying U.K. startup, U.S. company and China startup for 1 billion. Mr. John Doe is the CEO of the company." [
+"type": "ORG", "text": "U.K. startup",
+"type": "ORG", "text": "U.S. company",
+"type": "ORG", "text": "China startup",
+"type": "ORG", "text": "Apple",
+"type": "PER", "text": "John Doe",
+"type": "MISC", "text": "CEO",
+"type": "NUM", "text": "1 billion"
+Example 2 "Skai TV is a Greek free-to-air television network based in Piraeus.
+It is part of the Skai Group, one of the largest media groups in the country.
+It was relaunched in its present form on 1st of April 2006 in the Athens metropolitan area, and gradually spread its coverage nationwide.
+Besides digital terrestrial transmission, it is available on the subscription-based encrypted services of Nova and Cosmote TV.
+Skai TV is also a member of Digea, a consortium of private television networks introducing digital terrestrial transmission in Greece.
+At launch, Skai TV opted for dubbing all foreign language content into Greek, instead of using subtitles.
+This is very uncommon in Greece for anything except documentaries (using voiceover dubbing) and children’s programmes (using lip-synced dubbing), so after intense criticism the station switched to using subtitles for almost all foreign shows." [
+"type": "ORG", "text": "Skai TV",
+"type": "LOC", "text": "Greek",
+"type": "MISC", "text": "television network",
+"type": "LOC", "text": "Piraeus",
+Respond with a json format strictly.
+Table 14: The instruction for using LLMs for NER.
+C.4 Instruction for Document-level Relation Extraction
+We use the prompt in Table 15 with GPT-4 to conduct Document-level Relation Extraction.
 Instruction: Given a text and entities in the given text, your task is to extract document-level relation triples from the text.
@@
 (’Alexeni’, ’country’, ’Romania’)
-Table 10: The instruction for Document-level Relation Extraction
-A.4 Instruction for LLM validation
-We use the prompt below to instruct LLMs focus on one piece of information each time.
+Table 15: The instruction for DocRE.
+C.5 Instruction for LLM validation
+We use the prompt in Table 16 to instruct LLMs to focus on one piece of information each time.
 Instruction: Given a hypothesis and a reference, determine if the hypothesis is entailed by the reference.
@@
 Table 16: The instruction for LLM Validation
-A.5 Instruction for LLM-based Evaluation of Informativenss
+C.6 Instruction for LLM-based Evaluation of Informativeness
 We use the following prompt in Table 17 to instruct LLMs to generate the informative score.
@@
 0.228
-Table 12: The instruction for Document-level Relation Extraction
-Appendix B Case Study
-To provide a more intuitive and straightforward insight into the pipeline of EVA-Score, we provide a detailed case.
-From the case, we can see that EVA-Score provides not only a quantitative and objective information score but also provides a human-like explanation.
-Reference Summary:
-The mpp decreased the number of dopaminergic neurons by 40% , and increased the release of lactate dehydrogenase ( ldh ) into the culture medium .
-The tq significantly rescued dopaminergic neurons and decreased the release of ldh at the concentrations of 0.1 and 1 m
-Atomic Fact Chains
-[- The mpp decreased dopaminergic neurons, - The mpp decreased the number of dopaminergic neurons, - The mpp decreased the number of dopaminergic neurons by 40%]
-[- The mpp increased the release of lactate dehydrogenase ( ldh ), - The mpp increased the release of lactate dehydrogenase ( ldh ) into the culture medium]
-[- The tq significantly rescued dopaminergic neurons]
-[- The tq decreased the release of ldh, - The tq decreased the release of ldh at the concentrations, - The tq decreased the release of ldh at the concentrations of 0.1 and 1 m]
-Document-level Relations:
-tq causes Thymoquinone
-mpp leads to 1-methyl-4-phenylpyridinium
-Candidate Summary:
-The MPP significantly decreased the survival of dopaminergic neurons and increased the release of lysosomal hydrolase (LDH) into the culture medium.
-The neurotoxicity of MPP was attributed to its selective uptake by dopaminergic neurons through the dopamine transporter.
-Atomic Fact Chains
-[- The MPP significantly decreased dopaminergic neurons, -The MPP significantly decreased the survival of dopaminergic neurons ]
-[- The MPP increased the release of lysosomal hydrolase (LDH), - The MPP increased the release of lysosomal hydrolase (LDH) into the culture medium]
-[- The neurotoxicity was attributed to dopaminergic neurons, - The neurotoxicity of MPP was attributed to dopaminergic neurons, - The neurotoxicity of MPP was attributed to its selective uptake by dopaminergic neurons, - The neurotoxicity of MPP was attributed to dopaminergic neurons through the dopamine transporter, - The neurotoxicity of MPP was attributed to its selective uptake by dopaminergic neurons through the dopamine transporter.]
-Document-level Relations:
-mpp causes 1-methyl-4-phenylpyridinium
-neurotoxicity infects neuroprotective effect
-Matching Process
-- “The mpp decreased dopaminergic neurons” match “The MPP significantly decreased dopaminergic neurons”
-- “The mpp increased the release of lactate dehydrogenase ( ldh )” match “The mpp increased the release of lactate dehydrogenase ( ldh )”
-- “The mpp increased the release of lactate dehydrogenase ( ldh ) into the culture medium” match “The MPP increased the release of lysosomal hydrolase (LDH) into the culture medium”
-- “The mpp decreased dopaminergic neurons and increased the release of lactate dehydrogenase ( ldh )” match ‘The MPP significantly decreased dopaminergic neurons and increased the release of lysosomal hydrolase (LDH)”
-- “The mpp decreased dopaminergic neurons and increased the release of lactate dehydrogenase ( ldh ) into the culture medium” match “The MPP significantly decreased dopaminergic neurons and increased the release of lysosomal hydrolase (LDH) into the culture medium”
-- “mpp causes 1-methyl-4-phenylpyridinium” match “mpp leads to 1-methyl-4-phenylpyridinium”
-Table 13: Case Study For EVA-Score from Pubmed Dataset.
+Table 17: The instruction for prompting ChatGPT and GPT-4 for evaluation.